### v1

# SAMSA: Efficient Transformer for Many Data Modalities

Abstract.
The versatility of self-attention mechanism earned transformers great success in almost all data modalities, with limitations on the quadratic complexity and difficulty of training. Efficient transformers, on the other hand, often rely on clever data-modality-dependent construction to get over the quadratic complexity of transformers. This greatly hinders their applications on different data modalities, which is one of the pillars of contemporary foundational modeling. In this paper, we lay the groundwork for efficient foundational modeling by proposing SAMSA - SAMpling-Self-Attention, a context-aware linear complexity self-attention mechanism that works well on multiple data modalities. Our mechanism is based on a differentiable sampling without replacement method we discovered. This enables the self-attention module to attend to the most important token set, where the importance is defined by data. Moreover, as differentiability is not needed in inference, the sparse formulation of our method costs little time overhead, further lowering computational costs. In short, SAMSA achieved competitive or even SOTA results on many benchmarks, while being faster in inference, compared to other very specialized models. Against full self-attention, real inference time significantly decreases while performance ranges from negligible degradation to outperformance. We release our source code in the repository: https://github.com/HySonLab/SAMSA.
differentiable algorithm,efficient transformer,deep learning
Correspondence to Truong Son Hy at thy@uab.edu
ccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning
1. Introduction
Transformers (Vaswani et al., 2017) have been successfully applied to a wide variety of tasks on many data modalities, namely sequence (Vaswani et al., 2017; Radford et al., 2019; Devlin et al., 2019), vision (Dosovitskiy et al., 2021; Carion et al., 2020; Liu et al., 2023a), speech (Gulati et al., 2020; Gong et al., 2021; Koutini et al., 2022), time-series (Liu et al., 2022; Jin et al., 2024), point cloud (Lee et al., 2019; Zhao et al., 2021; Yang et al., 2019), and graph (Yun et al., 2019; Rampášek et al., 2022; Shirzad et al., 2023; Ngo et al., 2023). This can be attributed to the permutation equivariance of transformers, which provides a unique way of encoding structural information through positional encoding (absolute or relative positional enccodings). Ultimately, transformers become the building blocks of contemporary foundational modelings. While transformer can be the one architecture for all data modalities, quadratic complexity that hinders it from processing very long sequences. To mitigate the problem of quadratic complexity, based on the data modality they are working on (mostly sequence), efficient transformers (Beltagy et al., 2020; Ainslie et al., 2020; Zhu et al., 2021; Zaheer et al., 2020; Shirzad et al., 2023) have been developed. This is great for its original usage and some of the mentioned works have found their way to industrial applications. However, it has a major flaw: it only works for the intended data modality and is often hard to adapt to another data modality. For example, it is unreasonable to expect sliding window attention (Zhu et al., 2021) to work on data with non-rigid data structures like point clouds or graphs. Without efficient building blocks working across many data structures, future works toward AGI would be costly in both training and inference.
In this work, we develop towards data modality agnostic efficient transformer architecture to fill the literature gap. Any satisfying architecture should respect the permutation equivariance that full self-attention adheres to for versatility, be both sub-quadratic asymptotic computational cost for very large number of tokens, and be efficient in common cases (i.e. 512 - 1024 tokens) for practical usage. Inspired by the finite-yet-directed nature of human attention, we developed a parallelizable differentiable sampling without replacement algorithm that allows models to learn to choose what set of tokens they should attend to. This parallelization of our differentiable sampling algorithm enables our model to be trained efficiently while the without-replacement characteristic allows our model to maintain some functional expressivity. Since the result of one attention head is not dependent on the input of any other attention head, different attention heads can use different sampled sets of tokens. This consequentially enlarges the receptive field, which further increases the expressivity of our model.
To demonstrate the effectiveness of our method on many data modalities, we conduct extensive experiments on multilinear and non-multilinear data modalities. For multilinear data modalities, we experimented on sequence data modality via the long-range arena benchmark (Tay et al., 2021) featuring four classification tasks and one retrieval task on these datasets. For non-multilinear data modalities, we experimented on graph and point cloud data modalities: graph classification/regression via peptides dataset (Singh et al., 2015) from Long-Range Graph Benchmark (Dwivedi et al., 2022), object classification on ModelNet40 dataset (Wu et al., 2015), and part segmentation on ShapeNet dataset (Chang et al., 2015). Compared to specialized models with data-modality-specific inductive bias, we achieve competitive results while being much faster than them. Against full self-transformers, the performance degradation was negligible, and in some cases, unexpectedly outperformed full self-attention transformers. Aside from comparative experiments, we also conducted many ablation studies to understand: the efficiency-performance trade-off, regularization effect of sampling mechanisms, and training stability w.r.t number of sampled tokens and the sparse formulation.
In short, our contributions can be summarized as follows:
  • Parallelizable differentiable sampling without replacement method and its integration to self-attention mechanism,
  • Fast training time and negligible extra cost for inference compared to transformers with negligible performance loss,
  • Demonstrated effectiveness on many data modalities, competitive results even against specialized models.
2. Related works
Efficient Transformers.
The scalability of transformers is hindered by the quadratic complexity of self-attention modules. Hence, numerous works have delved into sparsity (Beltagy et al., 2020; Ainslie et al., 2020; Zaheer et al., 2020; Shirzad et al., 2023), low-rank projection (Wang et al., 2020), hashing-based method (Kitaev et al., 2020), and kernel-based methods (Choromanski et al., 2021; Katharopoulos et al., 2020) to sacrifice the expressivity of transformers for a sub-quadratic construction. Our review of contemporary efficient transformers is influenced by this survey by (Tay et al., 2022). Linformer (Wang et al., 2020) is based on the low-rank assumption of self-attention matrices, realized via projection matrices. As one of the earliest works in efficient transformers, it has flaws in both scalability and fixed sequence length. Subsequent work like Linear Transformer (Katharopoulos et al., 2020) redesigns the similarity kernel previously as a softmax of dot-product score (or any other kernels) to one made of a linear probability cage. The modification frees them from the recomputation of the multiplication of key-value matrices, making attention computational cost $O(N)$, where $N$ refers to the number of tokens. Without softmax non-linearity, the rank of those attention matrices is bounded by the number of dimensions of tokens, which loses much expressivity of transformer models. Methods like (Beltagy et al., 2020) utilize three choices of fixed-sparsity-patterns: sliding window, dilated sliding window, and global+sliding window. Similarly, (Zaheer et al., 2020) uses a combination of different sparse patterns at one time: random attention, window attention, and global attention. On graphs, (Shirzad et al., 2023) also combines different sparsity patterns but with a constraint on the random-generated one: it has to be an expander graph. The hard-coded sparsity effectively linearized the complexity of transformers and is empirically fast. However, intuitively, no pattern or finite set of patterns fits all problems; therefore, there is also a line of work dedicated to a differentiable sparsity. (Roy et al., 2021) drew a relationship from MIPS problem (Maximum Inner Product Search) and the Nearest-Neighbor Search algorithm, when both the query and key vectors are unit vectors. With their online k-means algorithm, their method attends each query vector to every key vector belonging to the same cluster, thus, bringing the complexity down to $O(n^{1.5})$, with $n$ is the number of tokens. Practically, however, their method is not fast as it needs a k-means construction and extensive specialized implementation for sparse operators. Another perspective on efficient transformers is making better implementation of vanilla transformers (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024). (Rabe and Staats, 2021) proposed a neat trick for attention mechanism: chunking. While the mechanism is simple, they have reduced the memory cost from quadratic to sublinear, effectively lifting the computation constraint of GPU memory. (Dao et al., 2022) both reduces the frequency of data transfer between HBM and SRAM within the GPU and derives a memory efficient backward pass. This allows quadratic attention scales to the sequence length of 64000 and beyond. (Dao, 2024) is an incremental work from (Dao et al., 2022) with better work partition and reduction in non-matmul FLOPS.
Continuous Relaxation
Discrete operations are much used in decision problems (classification, next-word prediction, plan, …). However, these operations by nature are not amenable to gradient-based optimization because that operation has gradients everywhere either zero or undefined. For example, the heaviside function for binary choice:
(1) $H(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$
has gradient undefined at $x=0$ and zero everywhere else. Since continuous relaxation of discrete operators and algorithms is a complex study and the sampling operator is the core of our efficient transformer formulation, we focus our literature review on differentiable sampling. To continuously relax the sampling operator, the previous literature was developed based on the relaxation of choose, sort, and top-k algorithms (Jang et al., 2017; Maddison et al., 2017; Kool et al., 2019; Xie et al., 2020; Blondel et al., 2020; Petersen et al., 2021). (Jang et al., 2017; Maddison et al., 2017) is the foundational work on parameterization of discrete categorical variable. They introduces the Gumbel-Softmax distribution that approximate categorical distribution exceptionally well. There are some differences between the two works: (Jang et al., 2017) introduces straight through estimator via application of path derivative (Bengio et al., 2013) and (Maddison et al., 2017) takes account of density in their relaxed objective function. From differentiable categorical variable, differentiable sampling with replacement can be obtained trivially by computing these stochastic categorical variable $k$-times, where $k$ is the number of samples. Differentiable sampling without replacement, however, is much trickier and is an active area of research. (Kool et al., 2019) is one of the earliest work in differentiable sampling without replacement. Their process of sampling without replacement, Gumbel Top-k, samples tokens sequentially with logit normalization, applied in generative text models as stochastic beam search. (Xie et al., 2020) sees the differentiable without replacement in another perspective: optimal transport, where loss function represents how far away is the current solution from the optimal one. (Blondel et al., 2020) is an efficient estimation of sort function differentiation through optimal transport via permutahedron projections. The projection decrease quadratic computational cost to $O(n\log{}n)$ forward pass and $O(n)$ backward pass. (Petersen et al., 2021) takes a pre-existing discrete parallelizable sorting/top-k algorithm (e.g. odd-even sorting network, bitonic sorting network, selection network, …) and relaxes all of the control flows for gradient estimation.
3. Background
3.1. Transformer Architecture
The self-attention mechanism is a core component of Transformer models to capture long-range dependencies. Let $\mathbf{X}=\left\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right\}$ $\in\mathbb{R}^{n\times d}$ represent a sequence of $n$ $d$-dimensional vectors. The matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ all in $\mathbb{R}^{n\times d_{h}}$ representing query, key, and value vectors are created using linearity:
$\mathbb{Q}$ $\;=\mathbb{X}W_{Q}+\mathbb{1}_{n}b_{q}^{\top},$
$\mathbb{K}$ $\;=\mathbb{X}W_{K}+\mathbb{1}_{n}b_{k}^{\top},$
$\mathbb{V}$ $\;=\mathbb{X}W_{V}+\mathbb{1}_{n}b_{v}^{\top}.$
Then, using the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ vectors, it creates another sequence by aggregating tokens satisfying certain criteria, expressed as follows:
(2) $\operatorname{Attn}(\mathbf{X})=\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V},$
where $\operatorname{SM}$ is the softmax nonlinearity. The output of the self-attention layer is propagated through a token-wise feed forward
(3) $\operatorname{FFN}(\mathbb{X})=\sigma(\mathbb{X}W_{1}+\mathbb{1}_{n}b_{1}^{\top})W_{2}+\mathbb{1}_{n}b_{2}^{\top},$
where $\sigma$ is an elementwise nonlinearity. The intrinsic problem of Transformer is the calculation of $\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{d_{h}}\right)$ in Eq.2, which calculate pair-wise attention score, leading to $O(n^{2})$ computation cost. Despite the computational complexity of transformers being quadratic, the memory complexity of transformers can be reduced to $O(n)$, shown in (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024). The computational cost of self-attention is further decreased through fewer memory operations between HBM and SRAM and better parallelization in (Dao et al., 2022; Dao, 2024). These improvements are compatible with our model and are integrated into our model via existing fused attention implementation.
3.2. Continuous Relaxation of Top-$k$ Sampling
A methodology to find the optimal sparse pattern for attention is too make such process learnable, i.e. differentiable and convergent. But evidently, top-$k$ sampling methods are not differentiable. A way to work around this is by relaxing the sampling function, represented by the sampling matrix, by an soft alternative. The criteria for choosing such a relaxation is such that there is some hyperparameter that can be tuned to approximate the hard function from the soft function with abitrary accuracy, for instance the softmax and the hardmax. The goal of this technique is to differentiablize functions without derivatives.
In the context of a top-$k$ sampling function $\mathcal{s}:\mathbb{R}^{n}\to\mathbb{R}^{k}$, these functions can be represented as a transformation $S(x)=\mathbb{P}x$ where $\mathbb{P}\in\{0,1\}^{k\times n}$ and each row of $\mathbb{P}$ contains one value 1. If $\mathbb{P}$ is fixed and independent of $x$, then $S(x)$ is differentiable and such methods are fixed sampling, or fixed sparsity patterns such as sliding window, random attention, this is not possible if we let $\mathbb{P}=\mathbb{P}(x)$. The common solution to this is approximating $\mathbb{P}$ by a softmax output, which is continuous. One can also use the soft approximation solely for the backward calculation, this is called the reparametrization trick. To examplify reparametrization trick, we present an optimization problem as a toy example.
Example 3.1.
Given an input vector $\mathbb{x}\in\mathbb{R}^{10}$ and a label $y\in\mathbb{R}^{5}$, minimize the weight sum of five values of $\mathbb{x}$ against $y$. In other words, solve the optimization problem
$\min_{\begin{subarray}{c}{i_{j},w,b}\\ {j\in[5]}\end{subarray}}\sum_{j=1}^{5}\left(x_{i_{j}}w+b-y_{j}\right)^{2}.$
To solve Example 3.1, our goal is to design a neural network which features sampling and obtains the three largest values. Let us call assume our network be the composition $f\circ g$ where $f$ be the simple neural network defined by $w$ and $b$, while $g$ is the sampling function selecting five most relevent inputs in $\mathbb{x}$. Assume that $f\circ g$ lies in the hypothesis space, which means that it can perfectly learn, i.e. find the hard threshold for three largest values of $\mathbb{x}$. In our design, the sampling function $g:\mathbb{R}^{10}\to\mathbb{R}^{5}$ should be
$g(\mathbb{x})=\mathbb{P(x)x}$
where $\mathbb{P}:\mathbb{R}^{10}\to\mathbb{R}^{5}$ is the sampling matrix that is value-aware. However, $g$ is clearly not differentiable, and, thus, cannot be trained in practice. To handle this, we create a backward soft proxy for $g$ called $g_{soft}$ where during training,
$\nabla g\triangleq\nabla g_{soft}$
but the forward process remains the same. This is the main concept of the reparametrization trick.
4. Method
4.1. Overview
Figure 1. Overview of our proposed model sampling-self-attention module SAMSA. The key-value vectors are selected via top-$k$ importance score computed using tokens’ latent. The Gumbel-Sigmoid reparameterization provides gradients to guide the optimization process towards most important key-value pairs of vectors (left). The sampled key-value vectors are then fed into Flash Attention to attend to the query vectors (right).
Our method is a composition of context-aware sampling without replacement method (Section 4.2) and the integration to self-attention mechanism (Section 4.3). Via reparameterization via Gumbel-Softmax, we devised a parallelizable differentiable sampling without replacement as an innovation from Gumbel-Top-K in existing literature (Kool et al., 2019) - a sequential method. Similar to Linformer (Wang et al., 2020) applying low-rank approximation to key and value vectors, we attend query vectors to the sampled key and value vectors. Inspired by the MoE architecture where each MLP chooses different subsets of tokens to transform (Zhou et al., 2022), our architecture assigns different sets of key-value pairs to different heads. This novel paradigm can be seen as mixture of attention experts, where each expert oversees different sets of tokens. The method is summarized in Algorithm 1.
Algorithm 1 Sampling-based Transformer Layer
0:  Input token matrix $\mathbf{X}$, sample size $k$
0:  Updated token matrix $\mathbf{X}$
1:  Transform token matrix to query, key, and value vectors
2:  $\widehat{\mathbb{X}}\leftarrow\operatorname{RMSNorm}(\mathbf{X})$
3:  $\mathbf{Q}\leftarrow\widehat{\mathbb{X}}W_{Q}+\mathbb{1}_{n}b_{q}^{\top}$
4:  $\mathbf{K}\leftarrow\widehat{\mathbb{X}}W_{K}+\mathbb{1}_{n}b_{k}^{\top}$
5:  $\mathbf{V}\leftarrow\widehat{\mathbb{X}}W_{V}+\mathbb{1}_{n}b_{v}^{\top}$
6:  Compute importance score
7:  $\mathbf{S}\leftarrow\operatorname{ImportanceScore}(\mathbf{X})+\operatorname{Gumbel}(0,1)$
8:  $\mathbb{P}\leftarrow\operatorname{SamplingMatrix}(\mathbb{S},k)$Refer to Algorithm 2}
9:  $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\leftarrow\mathbb{PK},\mathbb{PV}$
10:  Compute attention score
11:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FlashAttention}(\mathbf{Q},\tilde{\mathbf{K}},\tilde{\mathbf{V}})$
12:  Feed-Forward Network
13:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FFN}(\widehat{\mathbb{X}})$
14:  Return $\mathbf{X}$
4.2. Differentiable Top-$k$ Sampling
Sampling operator definition.
We formally define the operator that samples $k$ of $n$ real $d$-dimensional vectors as a linear function $\operatorname{S}(\mathbf{X},k):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{k\times d}=\mathbf{P}\mathbf{X}$. Here, $\mathbf{P}$ the sampling matrix is a row-wise concatenation of one-hot vectors defined as follows:
(4) $\mathbb{P}=\left[\begin{array}[]{ccccc}\operatorname{onehot}(x_{1})\\ \operatorname{onehot}(x_{2})\\ \vdots\\ \operatorname{onehot}(x_{k})\\ \end{array}\right].$
where $x_{1\dots k}$ indicates the one-hot vector represents the indices of the most important tokens.
Score-based Context-Aware Sampling.
Here, we define importance scores as the tokens’ likeliness of being picked. The importance scores of tokens are computed via a neural network: $\operatorname{NN}(\mathbf{X}):\mathbb{R}^{nm}\rightarrow\mathbb{R}^{n}$. While this method is simple, it conditions the importance score of each token on the entire set of tokens, allowing patterns with complex dependencies to emerge. For randomness, we use $\operatorname{Gumbel}$ distribution to include randomness to the computed importance scores, similar to (Jang et al., 2017). The Gumbel-Distribution PDF and importance score computation is given as follows:
$\operatorname{Gumbel}(0,1)(x)$ $=e^{-(x+e^{-x})},$
$\operatorname{ImportanceScore}(\mathbb{X})$ $=\operatorname{NN}(\mathbb{X}).$
With-replacement sampling.
Gumbel-Softmax (Jang et al., 2017) provides a differentiable stochastic choosing function via continuous relaxation of hardmax function. The sampling-with-replacement procedure via Gumbel-Softmax with reparameterization trick is given as follows:
$z$ $=\operatorname{ImportanceScore}(\mathbb{X}),$
$s(\mathbf{X},g)$ $=\mathbf{X}_{\operatorname{argmax}(z+g)},$
$\nabla s(\mathbf{X},g)$ $=\nabla\left(\sum_{i=1}^{n}\frac{\exp\left(\frac{z_{i}+g_{i}}{\tau}\right)}{\sum_{j=1}^{n}\exp\left(\frac{z_{j}+g_{j}}{\tau}\right)}\mathbf{X}_{i}\right),$
$S(\mathbf{X},k)$ $=\operatorname{Stack}_{i=1}^{k}(s(\mathbf{X},g)),$
where $S$, $s$, $\mathbf{X}$, $z$, $\tau$, $k$, and $g$ are the sampling function, the sampling one-vector function, the value matrix, the importance score, the Gumbel temperature hyperparameter, the number of sampled vectors, and Gumbel random variable, respectively. Since the selected vectors are independent, parallelization is trivial in sampling with replacement. This is the most important characteristic as our intended use-case of differentiable sampler in this work is as a component of deep transformer networks.
Without-replacement sampling.
While parallelization of sampling with replacement is so much desirable, the existence of duplicates harms expressivity. Therefore, we constructed a more desirable sampling method via reparameterization that rules out duplicates, is differentiable, and most importantly has great parallelizability. Our method estimates the gradient of hard sampling via a proxy soft-sampler. This soft-sampler selects top-2k important vectors. Then, for each of the top-k most important vectors, it compares against every of the k-less important vectors in k pairs with Gumbel-Sigmoid to generate gradients (a variant of Gumbel-Softmax where the number of vectors is two). While the top-k procedure is non-differentiable, putting vector pairs in probability bins makes those vectors compete for slots to be selected. As a result, through gradient descent, the more important vectors defined via data will have their score go toward up while the others go down. Mathematically, the method is described as follows:
$z$ $=\operatorname{ImportanceScore}(\mathbb{X})+g,$
$i$ $=\operatorname{ArgTopk}(z,2k),$
$p$ $=\operatorname{Sigmoid}\left(\frac{z_{i[k:]}\mathbb{1}_{k}^{\top}-\mathbb{1}_{k}z_{i[k:]}^{\top}}{\tau}\right),$
$S(\mathbf{X},k)$ $=\mathbf{X}_{i[:k]},$
$\nabla S(\mathbf{X},z,g)[j]$ $=\nabla\left(\sum_{g=1}^{k}\frac{\mathbb{X}[i[j]]*p[j,g]+\mathbb{X}[i[k]]*(1-p[j,g])}{k}\right)$
where $\mathbb{1}_{k}$ are the all-one vectors of length $k$.
Algorithm 2 Differentiable Sampling Without Replacement
0:  Importance score $\mathbb{S}\in\mathbb{R}^{n}$, number of sampled token $k$
0:  Sampling matrix $\mathbb{P}\in\mathbb{R}^{k\times n}$
1:  Select top-2k important score
2:  $\mathbb{I}\leftarrow\operatorname{ArgTopK}(\mathbf{S},2k)$
3:  Compute the proxy function
4:  $\mathbb{P}\leftarrow\mathbb{0}_{k\times n}$
5:  $\mathbb{S}\leftarrow\exp{(\mathbb{S})}$
6:  $\mathcal{S}\leftarrow\mathbb{I}_{:k}$
7:  $\mathcal{R}\leftarrow\mathbb{I}_{(k+1):2k}$
8:  $\sigma_{\mathcal{R}}\leftarrow 0$
9:  for $r$ in $\mathcal{R}$ do
10:     $\sigma_{\mathcal{R}}\leftarrow\sigma_{\mathcal{R}}+\mathbb{S}_{r}$
11:  end for
12:  for $i$ from $1$ to $k$ do
13:     $s\leftarrow\mathcal{S}_{i}$
14:     $\mathbb{P}_{i,s}\leftarrow\mathbb{S}_{s}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
15:     for $r$ in $\mathcal{R}$ do
16:        $\mathbb{P}_{i,r}\leftarrow\mathbb{S}_{r}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
17:     end for
18:  end for
19:  Return $(\operatorname{Onehot}(\mathbb{I}_{:k})-\mathbb{P}).\operatorname{detach()}-\mathbb{P}$
4.3. Integration to Self-Attention Mechanism
Learnable Support Vector Paddings.
In the previous section, we introduced our differentiable top-k sampling method, designed to enhance the efficiency of the self-attention mechanism. A straightforward integration of this method would involve allowing all query vectors to attend to a subset of sampled key and value vectors. However, this approach encounters a significant challenge: it fails when the number of tokens in a sequence is less than k, the predetermined number of sampled tokens. This limitation can cause the algorithm to break, especially when working with shorter sequences or in scenarios where the token count fluctuates.
To address this issue, we devised a solution involving the creation of 2k learnable key and value vectors, each associated with learnable importance scores. These vectors are sampled alongside the input tokens during the attention process. These additional vectors function similarly to support vectors in traditional kernel methods, providing additional decision boundaries within the self-attention mechanism. The introduction of these learnable support vectors not only mitigates the issue of insufficient tokens but also offers several other benefits. Firstly, they expand the learning capacity of transformers, allowing them to capture more nuanced patterns and relationships within the data. Secondly, these vectors stabilize the training process by ensuring that the model consistently attends to a fixed number of tokens, thereby maintaining the robustness and reliability of the attention mechanism across varying input sizes.
Mixture of Attention Experts.
While global attention mechanisms, such as the one we propose, offer significant advantages in terms of parallelizability and runtime efficiency, they are not without limitations. One key drawback is their restricted local receptive field, meaning that each token can only attend to the same subset of selected tokens within each layer. This limitation arises from the inherent design of global attention, where every token in the sequence attends to the same sampled tokens across all attention heads. However, upon closer examination, we realized that each attention head operates independently of the others. This independence implies that the outputs of the self-attention mechanism remain unchanged regardless of the token permutations applied within each attention head.
To overcome this limitation, we drew inspiration from the Mixture of Experts (MoE) literature, particularly the concept of expert-choosing-token (Zhou et al., 2022). Our method introduces a novel approach where each attention head selects a different subset of sampled tokens to attend to. This process can be executed in parallel, as the results of the h-sampling without replacement are independent of one another. By enabling each attention head to focus on a distinct subset of tokens, our transformer architecture effectively increases the local receptive field by a factor of h, where h represents the number of attention heads. This means that with 4 attention heads and a 25% sampling rate, our model can approximate a full self-attention pattern when necessary, significantly enhancing its ability to capture both global and local dependencies within the data.
Putting It All Together.
Our transformer architecture integrates several cutting-edge techniques to maximize performance and efficiency. We adopt the Pre-RMSNorm formulation (Chen et al., 2018; Zhang and Sennrich, 2019), which has been proven to be both faster and more conducive to a higher convergence rate during training. This formulation normalizes the input before applying the self-attention mechanism, leading to improved stability and efficiency. For the residual branch of our model, we employ a zero-initialized learnable scalar multiplier (Bachlechner et al., 2021), a technique known to enhance the training dynamics of deep networks. However, we introduce a slight modification: the learnable scalar multiplier is shared across all layers of the network, promoting consistency and reducing the number of parameters that need to be learned. For importance score computation, a multi-layer perceptron transforming tokens’ latent is sufficient because the latents are complexly transformed by multiple transformer layers before.
Modifications to integrate our sampling mechanism is described in earlier text. This integration allows our model to benefit from the efficiency and enhanced learning capacity provided by the differentiable top-k sampling and learnable support vector paddings. Finally, to align our model with the latest advancements in language technology, we incorporate flash attention (Dao et al., 2022; Dao, 2024). Flash attention is a highly optimized variant of self-attention that offers higher throughput and a lower memory footprint, making it ideal for large-scale models like ours. By combining these state-of-the-art techniques, our transformer architecture is not only capable of achieving top-tier performance across a variety of tasks but also does so with greater efficiency, ensuring it can be deployed effectively in real-world applications where computational resources and time are critical factors.
4.4. Complexity Analysis
Consider the input token $\mathbb{X}\in\mathbb{R}^{n\times d}$, query and key weights $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$, and $k$ sampled tokens, the computational cost for each 1-head transformer layer is as follows:
  • Linear transformations for $\mathbb{Q},\mathbb{K}$ and $\mathbb{V}$: $O(nd^{2})$,
  • Importance score of tokens for each attention head: $O(nd)$,
  • Selecting top-$k$ importance score: $O(n+k)$ forward pass, $O(n+k^{2})$ backward pass with the formulation in Section 4.2,
  • Attention from between $n$ queries and top-$k$ most important keys: $O(nkd_{h})$,
  • FFN network: $O(nd^{2})$.
Given these complexities and considering the long-ranged data structures which can implies $n\gg d$, the asymptotic computational complexity of one single SAMSA layer is $O(n)$. The asymptotic space complexity of one single SAMSA layer is also $O(n)$.
5. Experiments
Table 1. Performance/Runtime of SAMSA compared to the full self-attention (Vaswani et al., 2017) with different sampling sizes. Green and Red indicate better and worse performances/runtimes respectively.
SAMSA Full Attention
Sampling Size 128 128 256 256 512 512 N/A
Soft/Hard Sampling Soft Hard Soft Hard Soft Hard
Sequence ListOPS 41.53 41.13 41.89 40.57 41.33 40.88 37.11
Pathfinder 81.97 71.83 80.55 69.71 79.74 72.28 71.83
Image 48.64 48.24 48.73 47.38 48.24 46.75 42.94
Text 64.74 65.42 65.53 65.42 65.38 65.42 65.21
Retrieval - 82.05 - 81.89 - 81.81 79.14
Graph Pept-func 72.01 72.01 71.42 72.21 - 72.45 63.26
Pept-struct 0.2519 0.2458 - 0.2515 - 0.2468 0.2529
Point Cloud ModelNet40 91.11 90.91 91.72 90.95 91.35 91.39 91.07
ShapeNetPart - 85.74 - 85.87 - 85.62 85.1
Inference Speed (1024 tokens) 1.63×1.63$\times$1.63 × 2.17×2.17$\times$2.17 × 1.44×1.$4$ 1.62×1.62$\times$1.62 × 1.08×1.08$\times$1.08 × 1.44×1.$4$ 1$\times$
5.1. Experimental Design
Our SAMSA models utilize either hard or soft sampling techniques, determined by whether the reparameterization trick is employed or not. Soft sampling foregoes the reparameterization trick, instead operating with a very small Gumbel temperature (0.1), denoted as $g=g_{\text{soft}}$. This approach allows for a smoother, more probabilistic selection of tokens during training. Conversely, hard sampling employs the reparameterization trick with a standard Gumbel temperature (1.0), resulting in a more deterministic selection process. To rigorously evaluate the effectiveness of our approach, both in terms of performance and efficiency, we trained a total of 45 models. These models were benchmarked against full self-attention (specifically, Flash Attention v2 (Dao, 2024)) as well as other specialized state-of-the-art (SOTA) models across those three distinct data modalities:
  • Sequence (multilinear)
    • Long Range Arena (Tay et al., 2021)
  • Point Cloud (non-multilinear)
    • ModelNet40 (Wu et al., 2015)
    • ShapeNetPart (Chang et al., 2015)
  • Graph (non-multilinear)
    • Long Range Graph Benchmark (Peptides) (Dwivedi et al., 2022; Singh et al., 2015)
The selection of these datasets was intentional and rooted in a clear rationale. Our model, being a global attention model, was evaluated on datasets that highlight both the strengths and weaknesses typically associated with global attention mechanisms. Global attention models, while powerful in capturing long-range dependencies, are often criticized for their weakness in handling local structures. Several of the datasets we selected, such as 3 LRA datasets (Nangia and Bowman, 2018; Linsley et al., 2018; Krizhevsky and Hinton, 2009) and 2 LRGB datasets (Singh et al., 2015), are known to exhibit strong locality characteristics, making them ideal for testing this potential weakness. On the other hand, global attention models excel in tasks that require long-range modeling, a strength we tested using datasets that demand a global perspective, including 3 LRA datasets (Linsley et al., 2018; Maas et al., 2011; Radev et al., 2009), and 2 Point Cloud datasets (Wu et al., 2015; Chang et al., 2015).
Furthermore, the chosen data modalities—sequence, point cloud, and graph—are fundamentally different from each other, with minimal shared inductive biases. This diversity ensures that our model’s generalizability is rigorously tested, as it cannot rely on the exploitation of inherent structures unique to any one modality.
To ensure the validity and fairness of our comparisons with baseline models, we took several precautions to minimize bias. We avoided the use of data augmentation, which could introduce variability, and maintained nearly identical hyperparameters across different tasks. This consistent approach implies that we engaged in minimal hyperparameter tuning, allowing for a more accurate assessment of the SAMSA model’s inherent capabilities and its performance relative to other models under similar conditions. By controlling these variables, we aimed to present a clear and unbiased evaluation of SAMSA’s effectiveness across diverse data modalities and against a range of competitive benchmarks.
Table 2. (Long Range Arena) Accuracy on the full suite of long range arena (LRA) tasks.
Method ListOps Text Retrieval Image Pathfinder
BigBird (Zaheer et al., 2020) 36.05 64.02 59.29 40.83 74.87
Reformer (Kitaev et al., 2020) 37.27 56.10 53.40 38.07 68.50
Performer (Choromanski et al., 2021) 18.01 65.40 53.82 42.77 77.05
Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34
Luna-256 (Ma et al., 2021) 37.98 65.78 79.56 47.86 78.55
SAMSA-hard (Ours) 41.12 65.42 81.89 48.24 72.27
SAMSA-soft (Ours) 41.88 65.53 82.05 48.73 81.97
5.2. Sampling vs. Full-Self Attention
We evaluated the performance of our SAMSA model across different sampling rates, with detailed results presented in Table 1. The experiments were conducted on nine different datasets using six distinct SAMSA model configurations. The results consistently demonstrate that SAMSA not only outperforms the full self-attention mechanism in most cases but also exhibits superior efficiency in both training and inference phases. This performance boost is particularly notable given the reduced computational overhead associated with SAMSA.
Table 3. Experimental results on Peptides func and Peptides struct. We report the means and standard deviations of 4 random seeds.
Method
Peptides Struct
MAE
Peptides Func
AP
GCN (Kipf and Welling, 2017) 0.3496 ± 0.0013 0.5930 ± 0.0023
GINE (Xu et al., 2019) 0.6346 ± 0.0071 0.5498 ± 0.0079
GatedGCN (Bresson and Laurent, 2017) 0.3420 ± 0.0013 0.5864 ± 0.0077
GatedGCN + RWSE (Bresson and Laurent, 2017) 0.3357 ± 0.0006 0.6069 ± 0.0035
Drew-GCN + LapPE (Gutteridge et al., 2023) 0.2536 ± 0.0015 0.7150 ± 0.0044
GraphGPS + LapPE (Rampášek et al., 2023) 0.2500 ± 0.0005 0.6535 ± 0.0041
GRIT (Ma et al., 2023) 0.2460 ± 0.0012 0.6988 ± 0.0082
Graph VIT (He et al., 2023) 0.2449 ± 0.0016 0.6942 ± 0.0075
GraphMLP Mixer (He et al., 2023) 0.2475 ± 0.0015 0.6921 ± 0.0054
MPNN + VN + RWSE (Cai et al., 2023) 0.2529 ± 0.0009 0.6685 ± 0.0062
SAMSA-hard (ours) 0.2458 ± 0.0013 0.7221 ± 0.0042
SAMSA-soft (ours) 0.2519 ± 0.0011 0.7171 ± 0.0051
We hypothesize that SAMSA’s superior performance can be attributed to the regularization effect introduced by the sampling operator. During the early stages of training, the sampling mechanism tends to select tokens almost randomly, which is akin to the DropAttention approach (Zehui et al., 2019). This random selection acts as a form of regularization, preventing the model from overfitting by ensuring that attention is not overly focused on any particular subset of tokens. As training progresses and the sampler becomes more adept at identifying important tokens, the attention mechanism becomes more targeted, gradually diminishing the regularization effect. This dynamic behavior aligns with previously observed phenomena, where early-stage regularization accelerates the model’s convergence rate, leading to better overall performance (Liu et al., 2023b).
Regarding the trade-off between performance and efficiency, the data in Table 1 reveals a significant insight: SAMSA can achieve superior results with a sampling rate of just 256 tokens per head, outperforming both full self-attention and configurations with higher sampling rates. This finding strongly suggests that not all tokens contribute equally to the attention mechanism’s effectiveness. Instead, there exists an optimal or ”Goldilocks-zone” sampling rate where the regularization effect of sampling maximizes performance while simultaneously reducing inference time. This balance between efficiency and effectiveness is a key advantage of SAMSA, as it highlights the potential to streamline attention computation without sacrificing accuracy.
Table 4. Experimental results on ModelNet40 dataset (Wu et al., 2015) and part segmentation results on ShapeNetPart dataset (Chang et al., 2015) against point-cloud specialized baselines.
Method ModelNet40 ShapeNetPart
mAcc OA c. IoU i. IoU
PointNet (Charles et al., 2017) 86.2 89.2 80.4 83.7
Set Transformer (Lee et al., 2019) - 90.4 - -
PointNet++ (Qi et al., 2017) - 91.9 81.9 85.1
SpecGCN (Wang et al., 2018) - 92.1 - -
PointCNN (Li et al., 2018) 88.1 92.2 84.6 86.1
DGCNN (Wang et al., 2019) 90.2 92.2 82.3 85.1
PointWeb (Zhao et al., 2019) 89.4 92.3 - -
SpiderCNN (Xu et al., 2018) - 92.4 81.7 85.3
PointConv (Wu et al., 2019) - 92.5 82.8 85.7
Point2Sequence (Liu et al., 2019) 90.4 92.6 - 85.2
KPConv (Thomas et al., 2019) - 92.9 85.1 86.4
InterpCNN (Mao et al., 2019) - 93.0 84.0 86.3
Point Transformer (Zhao et al., 2021) 90.6 93.7 83.7 86.6
SAMSA-hard (Ours) 88.9 91.3 83.8 85.8
SAMSA-soft (Ours) 88.7 91.7 - -
5.3. Sampling vs. Specialized Models
A significant advantage of SAMSA is its applicability across a wide range of data modalities, making it a versatile tool in diverse fields. This versatility is evident in the performance comparisons shown in Table 2, Table 3, and Table 4, where SAMSA is evaluated against a variety of specialized methods in graph processing (using the Long Range Graph Benchmark) and point cloud analysis (using ModelNet40 and ShapeNetPart). Despite SAMSA being designed without any inherent inductive biases tailored to specific data types, it demonstrates a performance that is not only comparable to these specialized methods but also surpasses many graph-based models.
In the context of graph processing, SAMSA produced competitive results on both the Peptides Func and Struct datasets. Notably, it achieved an average precision (AP) of $72.21\%$, outperforming all other graph methods evaluated in these benchmarks. This highlights SAMSA’s ability to effectively capture and utilize the important features within graph data, even when competing against models specifically designed for such tasks.
For point cloud analysis, while SAMSA did not reach the same level of performance as in graph processing, primarily due to the absence of local information aggregation mechanisms which are crucial in this domain, it still delivered results that were on par with many models specifically optimized for point clouds. This demonstrates SAMSA’s robustness and adaptability, as it remains competitive even in scenarios where it faces inherent challenges. Overall, SAMSA’s ability to perform well across different data modalities underscores its potential as a foundational model, capable of replacing or complementing more specialized approaches in various applications.
6. Conclusion
In conclusion, the SAMSA (SAMpling-based Self-Attention) mechanism addresses the limitations of traditional transformers by reducing the quadratic complexity typically associated with self-attention. Unlike other efficient transformers that are often data-modality-specific, SAMSA provides a versatile, context-aware linear complexity solution that can be applied across various data modalities. By employing a differentiable sampling method to focus on the most important tokens, SAMSA achieves competitive or even state-of-the-art results on multiple benchmarks. Additionally, SAMSA’s sparse formulation during inference minimizes computational costs, offering faster inference times with little in performance, and in some cases, even outperforming full self-attention transformers.
Limitations.
Despite the promising results achieved with SAMSA, there are certain limitations on the biased gradients of hard sampling formulation, which makes our model hard to converge in more difficult problems (Pathfinder). The biased gradient issue in hard sampling methods is mostly from the misestimation of gradient of the non-chosen path. We believe that future work on more precise gradient estimation would yield inference-fast model with much better accuracy.
Appendix A Datasets
ModelNet40.
The ModelNet40 dataset (Wu et al., 2015) consists of 12,311 pre-aligned shapes divided into 40 classes, where the train and test sets consist of 9,843 instances and 2,468 instances respectively. ModelNet40 is the pioneer large-scale 3D CAD dataset. Unlike previous CAD datasets (Shilane et al., 2004), ModelNet40 is the pioneer large-scale dataset that is diverse in terms of both class and samples per class.
ShapeNetPart.
The ShapeNetPart dataset consists of 16,881 pre-aligned 3D shapes from 16 categories and is a part of a larger dataset: ShapeNetCore (51,300 3D models) (Chang et al., 2015). The 3D shapes in the ShapeNetPart dataset are annotated with 50 segmentation parts in total representing virtual real-world 3D semantic models. ShapeNet provides a diverse variety of shape annotations and corresponding shapes. The full ShapeNet dataset a is multitude containing upright and front orientation vectors, parts and keypoints, shape symmetries, but we only account for the part-segmenting task in our work.
Long Range Arena.
The Long Range Arena (LRA) (Tay et al., 2021) is a composition benchmark of 5 tasks: ListOps (Nangia and Bowman, 2018), ImDB review (Maas et al., 2011), ACL Anthology Network (Radev et al., 2009), Grayscaled CIFAR-10 (Krizhevsky and Hinton, 2009), and Pathfinder (Linsley et al., 2018). These five tasks are all classification and they feature very long sequences: ListOps (2,048 tokens), ImDB review (1,024 tokens), ACL Anthology Network (4,096 tokens), Grayscaled CIFAR-10 (1,024 tokens), and Pathfinder (1,024 tokens). All five tasks involves tackling long-ranged data structures and manifold categorizations, challenging networks’s generalization powers as well as memory and time efficiencies.
Long Range Graph Benchmark.
The Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) is a graph composition of five datasets. But in our work, we only consider two peptide molecular tasks: Peptides-struct and Peptides-func. The chemical structures and features of these peptides are represented as graph of around 155 nodes and diameter of around 57. These challenges are suitable for evaluating long range receptive fields of graph-based deep learning networks. Peptides-struct and Peptides-func are much similar, except the former is a graph regression task while the other is a graph classification task.
Appendix B Training Strategies and Data Preprocessing
We use the pytorch framework (Paszke et al., 2019) to conduct our empirical experiments.
Here is the list of hyperparameters:
  • Optimizer: AdamW (Loshchilov and Hutter, 2019)
  • Learning rate: 0.001
  • $\beta_{1},\beta_{2}$: 0.9, 0.999
  • Weight decay: 0.01
  • Learning scheduler: Cosine Annealing
  • Learning rate warm-up: Untuned Linear Warmup (Ma and Yarats, 2021)
  • Gradient Clipping: Max Gradient Norm (2.0)
Table 5. Training details.
Task Dataset Batch Size No. of Epochs
Point Cloud ModelNet40 32 600
ShapeNetPart 32 600
Sequence ListOPS 32 50
Text 32 50
Retrieval 32 50
Image 32 50
Pathfinder 32 50
Graph Peptides-struct 32 240
Peptides-func 32 160
Our training phase details can be visited in Table 5.
As for the input preprocessing, our works are as follows:
  • Long Range Arena: We simply use one-hot encoding vectors as inputs.
  • Long Range Graph Benchmark: We input the whole graph, with both node and edge as tokens.
  • ModelNet40: Following previous work (Charles et al., 2017), we use farthest point sampling to sample 1024 initial points.
  • ShapeNetPart: We operate on the initial 2,500 point-normal without any change to the dataset.
All of the training processes utilize Cross Entropy Loss, without label smoothing normalization.

### v2

@@
 SAMSA: Efficient Transformer for Many Data Modalities
-Abstract.
+Abstract
 The versatility of self-attention mechanism earned transformers great success in almost all data modalities, with limitations on the quadratic complexity and difficulty of training.
@@
 Against full self-attention, real inference time significantly decreases while performance ranges from negligible degradation to outperformance.
-We release our source code in the repository: https://github.com/HySonLab/SAMSA.
-differentiable algorithm,efficient transformer,deep learning
-Correspondence to Truong Son Hy at thy@uab.edu
-ccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning
-1. Introduction
+We release our source code in the repository: https://github.com/HySonLab/SAMSA
+$\dagger$\$$\dagger$\$footnotetext: Correspondence to thy@uab.edu
+1 Introduction
 Transformers (Vaswani et al., 2017) have been successfully applied to a wide variety of tasks on many data modalities, namely sequence (Vaswani et al., 2017; Radford et al., 2019; Devlin et al., 2019), vision (Dosovitskiy et al., 2021; Carion et al., 2020; Liu et al., 2023a), speech (Gulati et al., 2020; Gong et al., 2021; Koutini et al., 2022), time-series (Liu et al., 2022; Jin et al., 2024), point cloud (Lee et al., 2019; Zhao et al., 2021; Yang et al., 2019), and graph (Yun et al., 2019; Rampášek et al., 2022; Shirzad et al., 2023; Ngo et al., 2023).
@@
   • Demonstrated effectiveness on many data modalities, competitive results even against specialized models.
-2. Related works
+2 Related works
 Efficient Transformers.
@@
 The hard-coded sparsity effectively linearized the complexity of transformers and is empirically fast.
-However, intuitively, no pattern or finite set of patterns fits all problems; therefore, there is also a line of work dedicated to a differentiable sparsity.
+However, intuitively, no pattern or finite set of patterns fits all problems; therefore, there is also a line of work dedicated to a differentiable sparsity (Hy & Kondor, 2022; Nguyen et al., 2024).
 (Roy et al., 2021) drew a relationship from MIPS problem (Maximum Inner Product Search) and the Nearest-Neighbor Search algorithm, when both the query and key vectors are unit vectors.
@@
 Practically, however, their method is not fast as it needs a k-means construction and extensive specialized implementation for sparse operators.
-Another perspective on efficient transformers is making better implementation of vanilla transformers (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024).
-(Rabe and Staats, 2021) proposed a neat trick for attention mechanism: chunking.
+Another perspective on efficient transformers is making better implementation of vanilla transformers (Rabe & Staats, 2021; Dao et al., 2022; Dao, 2024).
+(Rabe & Staats, 2021) proposed a neat trick for attention mechanism: chunking.
 While the mechanism is simple, they have reduced the memory cost from quadratic to sublinear, effectively lifting the computation constraint of GPU memory.
@@
 For example, the heaviside function for binary choice:
-(1) $H(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$
+$H(x)=\begin{cases}0&\text{if }x<0\\ 1&\text{if }x\geq 0\end{cases}$ (1)
 has gradient undefined at $x=0$ and zero everywhere else.
@@
 (Blondel et al., 2020) is an efficient estimation of sort function differentiation through optimal transport via permutahedron projections.
-The projection decrease quadratic computational cost to $O(n\log{}n)$ forward pass and $O(n)$ backward pass.
+The projection decreases the quadratic computational cost to $O(n\log{}n)$ forward pass and $O(n)$ backward pass.
 (Petersen et al., 2021) takes a pre-existing discrete parallelizable sorting/top-k algorithm (e.g.
 odd-even sorting network, bitonic sorting network, selection network, …) and relaxes all of the control flows for gradient estimation.
-3. Background
-3.1. Transformer Architecture
-The self-attention mechanism is a core component of Transformer models to capture long-range dependencies.
-Let $\mathbf{X}=\left\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right\}$ $\in\mathbb{R}^{n\times d}$ represent a sequence of $n$ $d$-dimensional vectors.
-The matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ all in $\mathbb{R}^{n\times d_{h}}$ representing query, key, and value vectors are created using linearity:
-$\mathbb{Q}$ $\;=\mathbb{X}W_{Q}+\mathbb{1}_{n}b_{q}^{\top},$
-$\mathbb{K}$ $\;=\mathbb{X}W_{K}+\mathbb{1}_{n}b_{k}^{\top},$
-$\mathbb{V}$ $\;=\mathbb{X}W_{V}+\mathbb{1}_{n}b_{v}^{\top}.$
-Then, using the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ vectors, it creates another sequence by aggregating tokens satisfying certain criteria, expressed as follows:
-(2) $\operatorname{Attn}(\mathbf{X})=\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V},$
-where $\operatorname{SM}$ is the softmax nonlinearity. The output of the self-attention layer is propagated through a token-wise feed forward
-(3) $\operatorname{FFN}(\mathbb{X})=\sigma(\mathbb{X}W_{1}+\mathbb{1}_{n}b_{1}^{\top})W_{2}+\mathbb{1}_{n}b_{2}^{\top},$
-where $\sigma$ is an elementwise nonlinearity.
-The intrinsic problem of Transformer is the calculation of $\operatorname{SM}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{d_{h}}\right)$ in Eq.2, which calculate pair-wise attention score, leading to $O(n^{2})$ computation cost.
-Despite the computational complexity of transformers being quadratic, the memory complexity of transformers can be reduced to $O(n)$, shown in (Rabe and Staats, 2021; Dao et al., 2022; Dao, 2024).
-The computational cost of self-attention is further decreased through fewer memory operations between HBM and SRAM and better parallelization in (Dao et al., 2022; Dao, 2024).
-These improvements are compatible with our model and are integrated into our model via existing fused attention implementation.
-3.2. Continuous Relaxation of Top-$k$ Sampling
-A methodology to find the optimal sparse pattern for attention is too make such process learnable, i.e.
-differentiable and convergent.
-But evidently, top-$k$ sampling methods are not differentiable.
-A way to work around this is by relaxing the sampling function, represented by the sampling matrix, by an soft alternative.
-The criteria for choosing such a relaxation is such that there is some hyperparameter that can be tuned to approximate the hard function from the soft function with abitrary accuracy, for instance the softmax and the hardmax.
-The goal of this technique is to differentiablize functions without derivatives.
-In the context of a top-$k$ sampling function $\mathcal{s}:\mathbb{R}^{n}\to\mathbb{R}^{k}$, these functions can be represented as a transformation $S(x)=\mathbb{P}x$ where $\mathbb{P}\in\{0,1\}^{k\times n}$ and each row of $\mathbb{P}$ contains one value 1.
-If $\mathbb{P}$ is fixed and independent of $x$, then $S(x)$ is differentiable and such methods are fixed sampling, or fixed sparsity patterns such as sliding window, random attention, this is not possible if we let $\mathbb{P}=\mathbb{P}(x)$.
-The common solution to this is approximating $\mathbb{P}$ by a softmax output, which is continuous.
-One can also use the soft approximation solely for the backward calculation, this is called the reparametrization trick.
-To examplify reparametrization trick, we present an optimization problem as a toy example.
-Example 3.1.
-Given an input vector $\mathbb{x}\in\mathbb{R}^{10}$ and a label $y\in\mathbb{R}^{5}$, minimize the weight sum of five values of $\mathbb{x}$ against $y$.
-In other words, solve the optimization problem
-$\min_{\begin{subarray}{c}{i_{j},w,b}\\ {j\in[5]}\end{subarray}}\sum_{j=1}^{5}\left(x_{i_{j}}w+b-y_{j}\right)^{2}.$
-To solve Example 3.1, our goal is to design a neural network which features sampling and obtains the three largest values.
-Let us call assume our network be the composition $f\circ g$ where $f$ be the simple neural network defined by $w$ and $b$, while $g$ is the sampling function selecting five most relevent inputs in $\mathbb{x}$.
-Assume that $f\circ g$ lies in the hypothesis space, which means that it can perfectly learn, i.e.
-find the hard threshold for three largest values of $\mathbb{x}$.
-In our design, the sampling function $g:\mathbb{R}^{10}\to\mathbb{R}^{5}$ should be
-$g(\mathbb{x})=\mathbb{P(x)x}$
-where $\mathbb{P}:\mathbb{R}^{10}\to\mathbb{R}^{5}$ is the sampling matrix that is value-aware.
-However, $g$ is clearly not differentiable, and, thus, cannot be trained in practice.
-To handle this, we create a backward soft proxy for $g$ called $g_{soft}$ where during training,
-$\nabla g\triangleq\nabla g_{soft}$
-but the forward process remains the same. This is the main concept of the reparametrization trick.
-4. Method
-4.1. Overview
-Figure 1.
-Overview of our proposed model sampling-self-attention module SAMSA.
+3 Background
+3.1 Self-Attention Mechanism
+Vanilla Self Attention.
+The self-attention mechanism is a core component of Transformer models for long-range dependencies.
+Let $\mathbf{X}=\left\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right\}$ $\in\mathbb{R}^{n\times d}$ represent a set of tokens of $n$ $d$-dimensional vectors.
+The matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ all in $\mathbb{R}^{n\times d_{h}}$ representing query, key, and value vectors are created via learnable linear transformation of the input token matrix $\mathbf{X}$:
+$\mathbb{Q}$ $\;=\mathbb{X}W_{Q}+\mathbf{1}_{n}b_{q}^{\top},$
+$\mathbb{K}$ $\;=\mathbb{X}W_{K}+\mathbf{1}_{n}b_{k}^{\top},$
+$\mathbb{V}$ $\;=\mathbb{X}W_{V}+\mathbf{1}_{n}b_{v}^{\top}.$
+Then, using the query, key, and value vectors, it creates another set of tokens by aggregating those value vectors based on the dependencies defined through the attention map constructed by query and value vectors:
+$\operatorname{Attention}(\mathbf{X})=\operatorname{Softmax}\left(\frac{\mathbb{Q}\mathbb{K}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V},$ (2)
+This mechanism allows information aggregation of very far tokens and effectively captures complex long-range relationships.
+Efficient Quadratic Attention.
+Computing the given $\operatorname{Attention}$ function naively yields a quadratic memory cost due to the need to store attention maps for backpropagation.
+However, its memory complexity can be reduced into linear complexity by checkpointing (Rabe & Staats, 2021) or algebraic manipulation of derivative function (Dao et al., 2022).
+Via fused kernel (Dao et al., 2022; Dao, 2024), the computational cost of self-attention is further decreased through fewer memory operations between HBM and SRAM.
+This allows better parallelization, making quadratic attention scales to the length of 64k and beyond.
+3.2 Continuous Relaxation of Top-k Sampling
+Top-k Sampling.
+The top-k sampling operation selects the best set of $k$ objects out of others under a criterion, deterministic or stochastic.
+The criterion function is a permutation invariant function that maps a set of $n$ vectors to a real number.
+The following equation describes a mathematical formulation of top-k sampling:
+$\operatorname{Sample}(X,k)=\operatorname{argmax}_{S}^{C(X,k)}\text{criterion}(S),$ (3)
+where $X$ is the set of vector to sample from, $k$ is the number of sampled vectors needing to be sampled, and $C(N,k)$ is the set of all k-combinations of set $N$.
+Since the $\operatorname{argmax}$ operator has no continuity, the derivative of the sampled vectors with respect to the criterion is either zero or undefined, i.e., not useful.
+Therefore, it is hard to integrate into neural networks.
+Early work (Papernot & McDaniel, 2018) had to resort to a two-stage optimization.
+This circumvention, however, introduces additional complexity and worsens performance via non-aligned objectives.
+Straight Through Estimator.
+Before going deeper into the differentiable top-k literature, it is crucial to explain a much simpler case: $k=1$.
+To choose one vector from many, (Jang et al., 2017) provides a differentiable approximation of the categorical distribution via Softmax with Gumbel noise added on logits:
+$\operatorname{GumbelSoftmax}(\pi)[i]\triangleq\frac{\exp\left((\pi_{i}+g_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left((\pi_{j}+g_{j})/\tau\right)},$
+where $g_{i}$ are independent and identically samples from Gumbel(0,1) distribution, $\pi_{i}$ are the unnormalized logits, and $\tau$ is the temperature parameter that controls the smoothness of the distribution.
+By adjusting $\tau$, the Gumbel-Softmax can transition between a one-hot categorical distribution and a softer distribution, enabling gradient-based optimization.
+The straight-through estimator, however, uses the above soft-categorical function as a differentiable proxy for the hard-categorical function obtained by $\operatorname{argmax}$:
+$\operatorname{ST-GumbelSoftmax}(\pi)[i]$$\triangleq\begin{cases}1,&\text{if }i=\operatorname{argmax}(\pi),\\ 0,&\text{otherwise},\end{cases}$
+$\partial{\operatorname{ST-GumbelSoftmax}}/\partial{\pi}$ $\triangleq\partial{\operatorname{GumbelSoftmax}}/\partial{\pi}.$
+A naive way to extend this to differentiable top-k (Kool et al., 2019) is to use Gumbel-Softmax to sample from the unsampled vectors $k$ times sequentially, named Gumbel-Topk.
+This is not parallelizable and has unreliable gradients; therefore, it cannot be used inside repeating deep learning modules, e.g.
+transformer layers.
+4 Method
+4.1 Overview
+Figure 1: Overview of our proposed model sampling-self-attention module SAMSA.
 The key-value vectors are selected via top-k importance score computed using tokens’ latent.
@@
 0:  Updated token matrix $\mathbf{X}$
-1:  Transform token matrix to query, key, and value vectors
+1:  Transform tokens to query and key-value matrices (Pre-RMSNorm)
 2:  $\widehat{\mathbb{X}}\leftarrow\operatorname{RMSNorm}(\mathbf{X})$
-3:  $\mathbf{Q}\leftarrow\widehat{\mathbb{X}}W_{Q}+\mathbb{1}_{n}b_{q}^{\top}$
-4:  $\mathbf{K}\leftarrow\widehat{\mathbb{X}}W_{K}+\mathbb{1}_{n}b_{k}^{\top}$
-5:  $\mathbf{V}\leftarrow\widehat{\mathbb{X}}W_{V}+\mathbb{1}_{n}b_{v}^{\top}$
-6:  Compute importance score
-7:  $\mathbf{S}\leftarrow\operatorname{ImportanceScore}(\mathbf{X})+\operatorname{Gumbel}(0,1)$
-8:  $\mathbb{P}\leftarrow\operatorname{SamplingMatrix}(\mathbb{S},k)$Refer to Algorithm 2}
-9:  $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\leftarrow\mathbb{PK},\mathbb{PV}$
-10:  Compute attention score
-11:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FlashAttention}(\mathbf{Q},\tilde{\mathbf{K}},\tilde{\mathbf{V}})$
-12:  Feed-Forward Network
-13:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FFN}(\widehat{\mathbb{X}})$
-14:  Return $\mathbf{X}$
-4.2. Differentiable Top-$k$ Sampling
-Sampling operator definition.
-We formally define the operator that samples $k$ of $n$ real $d$-dimensional vectors as a linear function $\operatorname{S}(\mathbf{X},k):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{k\times d}=\mathbf{P}\mathbf{X}$.
-Here, $\mathbf{P}$ the sampling matrix is a row-wise concatenation of one-hot vectors defined as follows:
-(4) $\mathbb{P}=\left[\begin{array}[]{ccccc}\operatorname{onehot}(x_{1})\\ \operatorname{onehot}(x_{2})\\ \vdots\\ \operatorname{onehot}(x_{k})\\ \end{array}\right].$
-where $x_{1\dots k}$ indicates the one-hot vector represents the indices of the most important tokens.
-Score-based Context-Aware Sampling.
-Here, we define importance scores as the tokens’ likeliness of being picked.
-The importance scores of tokens are computed via a neural network: $\operatorname{NN}(\mathbf{X}):\mathbb{R}^{nm}\rightarrow\mathbb{R}^{n}$.
-While this method is simple, it conditions the importance score of each token on the entire set of tokens, allowing patterns with complex dependencies to emerge.
-For randomness, we use $\operatorname{Gumbel}$ distribution to include randomness to the computed importance scores, similar to (Jang et al., 2017).
-The Gumbel-Distribution PDF and importance score computation is given as follows:
-$\operatorname{Gumbel}(0,1)(x)$ $=e^{-(x+e^{-x})},$
-$\operatorname{ImportanceScore}(\mathbb{X})$ $=\operatorname{NN}(\mathbb{X}).$
+3:  $\mathbf{Q}\leftarrow\widehat{\mathbb{X}}\mathbf{W_{q}}+\mathbb{1}_{n}b_{q}^{\top}$
+4:  $\mathbf{P}\leftarrow\widehat{\mathbb{X}}\mathbf{W_{kv}}+\mathbb{1}_{n}b_{kv}^{\top}$
+5:  Compute multi-head importance scores
+6:  $\mathbf{Z}\leftarrow\operatorname{MultiHeadImportanceScore}(\mathbf{X})$
+7:  $\mathbf{Z}\leftarrow\mathbf{Z}+\operatorname{Gumbel}(0,1).\operatorname{sample()}$
+8:  Concatenate with 2k support vector paddings
+9:  $\mathbf{P}\leftarrow\operatorname{Concatenate}(\mathbf{P},\mathbf{P_{\text{supp}}})$
+10:  $\mathbf{Z}\leftarrow\operatorname{Concatenate}(\mathbf{Z},\mathbf{Z}_{\text{supp}})$
+11:  Sample sets of key and value vectors
+12:  $\tilde{\mathbf{P}}\leftarrow\operatorname{MultiHeadSample}(\mathbf{Z},\mathbf{P},k)$
+13:  $\tilde{\mathbf{K}},\tilde{\mathbf{V}}\leftarrow\operatorname{Split}(\tilde{\mathbf{P}})$
+14:  Attend and transform tokens
+15:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FlashAttention}(\mathbf{Q},\tilde{\mathbf{K}},\tilde{\mathbf{V}})$
+16:  $\widehat{\mathbb{X}}\leftarrow\operatorname{RMSNorm}(\mathbf{X})$
+17:  $\mathbf{X}\leftarrow\mathbf{X}+\operatorname{FFN}(\widehat{\mathbb{X}})$
+18:  Return $\mathbf{X}$
+4.2 Parallelizable Differentiable Top-k Sampling
+To be a viable key-value sampler for transformers, that method needs both great parallelizability and reliable gradients for efficient model training.
+Here, we introduce the two methods that hold the needed properties: trivial differentiable sampling with replacement and our proposed differentiable without replacement method.
 With-Replacement Sampling.
-Gumbel-Softmax (Jang et al., 2017) provides a differentiable stochastic choosing function via continuous relaxation of hardmax function.
-The sampling-with-replacement procedure via Gumbel-Softmax with reparameterization trick is given as follows:
-$z$ $=\operatorname{ImportanceScore}(\mathbb{X}),$
-$s(\mathbf{X},g)$ $=\mathbf{X}_{\operatorname{argmax}(z+g)},$
-$\nabla s(\mathbf{X},g)$ $=\nabla\left(\sum_{i=1}^{n}\frac{\exp\left(\frac{z_{i}+g_{i}}{\tau}\right)}{\sum_{j=1}^{n}\exp\left(\frac{z_{j}+g_{j}}{\tau}\right)}\mathbf{X}_{i}\right),$
-$S(\mathbf{X},k)$ $=\operatorname{Stack}_{i=1}^{k}(s(\mathbf{X},g)),$
-where $S$, $s$, $\mathbf{X}$, $z$, $\tau$, $k$, and $g$ are the sampling function, the sampling one-vector function, the value matrix, the importance score, the Gumbel temperature hyperparameter, the number of sampled vectors, and Gumbel random variable, respectively.
-Since the selected vectors are independent, parallelization is trivial in sampling with replacement.
-This is the most important characteristic as our intended use-case of differentiable sampler in this work is as a component of deep transformer networks.
+We define $z$ as the importance score of vectors, representing the vectors’ likeliness of being picked.
+The importance scores of vectors are computed via a neural network: $\operatorname{ImportanceScore}(\mathbf{X}):\mathbb{R}^{n\times m}\rightarrow\mathbb{R}^{n}$:
+$z\triangleq\operatorname{ImportanceScore}(\mathbb{X}).$
+This simple method conditions the importance score of each token on the entire set of tokens, allowing patterns with complex dependencies to emerge.
+If we allow the set of sampled vectors to contain duplicates, Gumbel-Top-k algorithm can be parallelized simply by computing everything concurrently as follows:
+$\operatorname{Sample}_{1}(z,\mathbf{X},k)$ $\triangleq\operatorname{Concatenate}_{i=1}^{k}(\operatorname{ST-GumbelSoftmax}(z)^{\top}\mathbf{X}).$
 Without-Replacement Sampling.
-While parallelization of sampling with replacement is so much desirable, the existence of duplicates harms expressivity.
-Therefore, we constructed a more desirable sampling method via reparameterization that rules out duplicates, is differentiable, and most importantly has great parallelizability.
-Our method estimates the gradient of hard sampling via a proxy soft-sampler.
-This soft-sampler selects top-2k important vectors.
-Then, for each of the top-k most important vectors, it compares against every of the k-less important vectors in k pairs with Gumbel-Sigmoid to generate gradients (a variant of Gumbel-Softmax where the number of vectors is two).
-While the top-k procedure is non-differentiable, putting vector pairs in probability bins makes those vectors compete for slots to be selected.
-As a result, through gradient descent, the more important vectors defined via data will have their score go toward up while the others go down.
-Mathematically, the method is described as follows:
-$z$ $=\operatorname{ImportanceScore}(\mathbb{X})+g,$
-$i$ $=\operatorname{ArgTopk}(z,2k),$
-$p$ $=\operatorname{Sigmoid}\left(\frac{z_{i[k:]}\mathbb{1}_{k}^{\top}-\mathbb{1}_{k}z_{i[k:]}^{\top}}{\tau}\right),$
-$S(\mathbf{X},k)$ $=\mathbf{X}_{i[:k]},$
-$\nabla S(\mathbf{X},z,g)[j]$ $=\nabla\left(\sum_{g=1}^{k}\frac{\mathbb{X}[i[j]]*p[j,g]+\mathbb{X}[i[k]]*(1-p[j,g])}{k}\right)$
-where $\mathbb{1}_{k}$ are the all-one vectors of length $k$.
-Algorithm 2 Differentiable Sampling Without Replacement
-0:  Importance score $\mathbb{S}\in\mathbb{R}^{n}$, number of sampled token $k$
-0:  Sampling matrix $\mathbb{P}\in\mathbb{R}^{k\times n}$
-1:  Select top-2k important score
-2:  $\mathbb{I}\leftarrow\operatorname{ArgTopK}(\mathbf{S},2k)$
-3:  Compute the proxy function
-4:  $\mathbb{P}\leftarrow\mathbb{0}_{k\times n}$
-5:  $\mathbb{S}\leftarrow\exp{(\mathbb{S})}$
-6:  $\mathcal{S}\leftarrow\mathbb{I}_{:k}$
-7:  $\mathcal{R}\leftarrow\mathbb{I}_{(k+1):2k}$
-8:  $\sigma_{\mathcal{R}}\leftarrow 0$
-9:  for $r$ in $\mathcal{R}$ do
-10:     $\sigma_{\mathcal{R}}\leftarrow\sigma_{\mathcal{R}}+\mathbb{S}_{r}$
-11:  end for
-12:  for $i$ from $1$ to $k$ do
-13:     $s\leftarrow\mathcal{S}_{i}$
-14:     $\mathbb{P}_{i,s}\leftarrow\mathbb{S}_{s}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
-15:     for $r$ in $\mathcal{R}$ do
-16:        $\mathbb{P}_{i,r}\leftarrow\mathbb{S}_{r}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
-17:     end for
-18:  end for
-19:  Return $(\operatorname{Onehot}(\mathbb{I}_{:k})-\mathbb{P}).\operatorname{detach()}-\mathbb{P}$
-4.3. Integration to Self-Attention Mechanism
+The weakness of sampling with replacement is that it creates duplicates.
+These duplicates waste computational resources on the same pieces of information, which harms the expressivity of transformers relying on sampling.
+This proposed method of without-replacement sampling relies on a novel proxy soft gradient estimator we devised.
+To begin with, we started with a hypothetical inefficient but parallelizable method: select one set of sampled vectors from sets of all sets of sampled vectors using Gumbel-Softmax (this is the definition in Equation 3).
+This is equivalent to a continuous relaxation of the discrete brute-force combinatorial optimization algorithm.
+By defining the likeliness of choosing one set (i.e.
+the importance score of sets of vectors) as the sum of the importance scores of its elements, we have:
+$z^{*}$ $\triangleq\operatorname{Concatenate}_{\mathbf{S}}^{C(\mathbb{N}^{\leq n},k)}(\sum_{i}^{\mathbf{S}}z_{i}),$
+$\mathbf{X}^{*}$ $\triangleq\operatorname{Concatenate}_{\mathbf{S}}^{C(\mathbb{N}^{\leq n},k)}(\operatorname{vec}(\mathbf{X}_{\mathbf{S}})^{\top}),$
+$\operatorname{Sample}_{2}(z,\mathbf{X},k)$ $\triangleq\operatorname{ST-GumbelSoftmax}(z^{*})^{\top}\mathbf{X}^{*},$
+where $C(N,k)$ is the set of all k-combinations of set $N$, $\mathbb{N}^{\leq n}$ is the set of all natural numbers less than or equal to $n$, $\operatorname{vec}$ is the matrix vectorization function obtained by stacking the columns of the input matrix on top of one another, $\mathbf{X}_{\mathbf{S}}$ is the matrix of the row vectors with indices in $\mathbf{S}$ of $\mathbf{X}$.
+This definition of the importance scores of sets has a clear advantage: not needing to enumerate over the entire $C(n,k)$ at inference.
+Since using all possible combinations to generate one gradient vector for a single optimization step is overkill, an immediate improvement is local search - a sequential algorithm.
+Intuitively, this allows us to put the optimization steps of local search and gradient descent in the same for loop.
+For local search, we define the current solution as the set with the highest importance score.
+The locality of an index set $S$ is then described as the collection of index sets, each differing from $S$ by exactly one element:
+$\operatorname{Current}$ $\triangleq\operatorname{argmax}(z^{*})=\operatorname{ArgTopK}(z)$
+$\operatorname{Local}(\mathbf{S})$$\triangleq\{x:x\in C(\mathbf{X},k)\text{ and }\left|x-\mathbf{S}\right|=1\}$
+Using local search as the base discrete combinatorial optimization algorithm, we compare the current solution against one of its localities, using Gumbel-Softmax reparameterization.
+The average result from comparing $k(n-k)$ pairs of solutions is given as follows:
+$i$ $\triangleq\operatorname{ArgTopK}(z,k),$
+$j$ $\triangleq\operatorname{ArgTopK}(-z,n-k),$
+$p$ $\triangleq J_{k,n-k},$
+$\partial p/\partial z$ $\triangleq\partial\left(\operatorname{ST-GumbelSigmoid}\left(z_{i}\mathbb{1}_{n-k}^{\top}-\mathbb{1}_{k}z_{j}^{\top}\right)\right)/\partial z,$
+$\operatorname{Sample}(z,\mathbf{X},k)[m]$ $\triangleq\sum_{v=1}^{n-k}\frac{\mathbb{X}[i_{m}]*p[m,v]+\mathbb{X}[j_{v}]*(1-p[m,v])}{n-k},$
+where $J_{m,n}$ is all-ones matrix with $m$ rows and $n$ columns, $\mathbb{1}_{k}$ are the all-one vectors of length $k$.
+The Gumbel Sigmoid function is the Gumbel Softmax when the number of categories is equal to two (details in Appendix D), similar to the relationship between Sigmoid and Softmax functions.
+To further cut down the optimization cost, we select $i$ as top-k most important and $j$ as top-k second most important index vectors.
+4.3 Graph Modeling
+The complexity of the graph data structure is needed for the claim of working on many data modalities of ours.
+Unfortunately, existing literature deemed self-attention alone insufficient for graph and they created very complicated methods: e.g.
+combination between Message Passing Neural Network (MPNN) and self-attention layers (Rampášek et al., 2022; Ying et al., 2021; Yun et al., 2019).
+This makes it hard to adapt to other data modalities.
+We identify the reason transformers need so many modifications just to be barely working on graphs is the inability to process both the nodes and edges information simultaneously.
+Therefore, we propose Graph-Bridge, a simpler approach treating both nodes and edges as tokens, which can be applied to any self-attention network without complicated constructions.
+Formulation.
+Given a graph $G=(V,E)$ where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to learn a mapping $\mathcal{F}:G\rightarrow\mathbb{R}^{d}$ that captures the structural properties of $G$ in a $d$-dimensional latent space.
+Each node $v_{i}\in V$ is associated with a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{f}$, and each edge $e_{ij}=(v_{i},v_{j})\in E$ is associated with a feature vector $\mathbf{e}_{ij}\in\mathbb{R}^{g}$.
+Positional Encoding.
+To integrate node and edge information, we introduce a novel graph positional encoding (PE) mechanism.
+Let $\mathbf{p}_{i}\in\mathbb{R}^{d}$ denote the positional encoding of node $v_{i}$.
+We define $\mathbf{p}_{i}$ as a random vector sampled from a normal distribution:
+$\mathbf{p}_{i}\sim\mathcal{N}(0,\operatorname{diag}(\sigma^{2}))+\phi(x_{i}),$
+where $\sigma^{2}$ is a learnable variance $d$-dimensional vector and $\phi:\mathbb{R}^{f}\rightarrow\mathbb{R}^{d}$ is a multi-layer perceptrons (MLPs) to inject node feature into positional encodings.
+For each edge $e_{ij}$, its positional encoding $\mathbf{p}_{ij}$ is computed as:
+$\mathbf{p}_{ij}=\mathbf{p}_{i}-\mathbf{p}_{j}.$
+While summation is usually used to combine positional encodings, in this case, we use subtraction for edge encoding to preserve the directional information of graph edges.
+Node and Edge Transformation.
+To map node and edge representations into a common latent space, we use two separate MLPs, $\phi_{V}:\mathbb{R}^{d}\times\mathbb{R}^{f}\rightarrow\mathbb{R}^{d}$ and $\phi_{E}:\mathbb{R}^{d}\times\mathbb{R}^{g}\rightarrow\mathbb{R}^{d}$, defined as:
+$\mathbf{h}_{i}$ $=\phi_{V}(\mathbf{p}_{i},\mathbf{x}_{i}),$
+$\mathbf{h}_{ij}$ $=\phi_{E}(\mathbf{p}_{ij},\mathbf{e}_{ij}),$
+where $\mathbf{h}_{i}$ and $\mathbf{h}_{ij}$ are the transformed representations of node $v_{i}$ and edge $e_{ij}$, respectively.
+These features are then processed by subsequent transformer layers, allowing complex relationships between nodes and edges to be recognized.
+Compare to other Graph-PE.
+This method is simple and efficient to compute: requiring only random feature generation, gathering operation on the randomly generated node PE for edges, and MLP features.
+This is different to existing graph positional encodings like Random-Walk PE (Ma et al., 2023a) (hard to parallelize) and Laplacian PE (Rampášek et al., 2022) (inherently computationally expensive).
+While being different, how this method works is surprisingly similar to random-walk PE: each self-attention propagates signals through edge tokens, i.e.
+we have merged the sequential process of random-walk and multi-layer neural network altogether.
+4.4 Integration to Self-Attention Mechanism
+While a straightforward integration of our sampling method into the self-attention mechanism is functional, we can achieve significant improvements. The following sections outline our enhancements.
+Mixture of Attention Experts.
+While global attention mechanisms are highly parallelizable and efficient, they limit each token’s local receptive field since all tokens attend to the same set of selected tokens.
+Each attention head operates independently, making self-attention outputs invariant to token permutations within each head.
+Inspired by MoE (Mixture of Experts) (Zhou et al., 2022), we propose a method where each attention head selects a distinct subset of sampled tokens.
+This process is parallelizable since the $h$-sampling processes are independent.
+Formally, the multi-head importance score $\mathbf{Z}$ is computed by:
+$\mathbf{Z}\triangleq\operatorname{MultiHeadImportanceScore}(\mathbb{X}).$
+And the multi-head sampling function $\operatorname{MultiHeadSample}$ as follows:
+$\mathbf{\tilde{P}}$ $\triangleq\operatorname{MultiHeadSample}(\mathbf{Z},\mathbf{P},k)$
+$=\operatorname{Stack}_{i=1}^{h}(\operatorname{Sample}(\mathbf{Z}_{\dots,i},\mathbf{P},k)),$
+where $\mathbf{Z}_{\dots,i}$ is the $i^{\text{th}}$ column vector of $\mathbf{Z}$ matrix.
+By incorporating this mixture of attention experts, our sampling transformer significantly enhances its local receptive field by a factor of $h$.
+For instance, a transformer with four attention heads and a 25% sampling rate can learn a full self-attention pattern if necessary.
 Learnable Support Vector Paddings.
-In the previous section, we introduced our differentiable top-k sampling method, designed to enhance the efficiency of the self-attention mechanism.
-A straightforward integration of this method would involve allowing all query vectors to attend to a subset of sampled key and value vectors.
-However, this approach encounters a significant challenge: it fails when the number of tokens in a sequence is less than k, the predetermined number of sampled tokens.
-This limitation can cause the algorithm to break, especially when working with shorter sequences or in scenarios where the token count fluctuates.
-To address this issue, we devised a solution involving the creation of 2k learnable key and value vectors, each associated with learnable importance scores.
-These vectors are sampled alongside the input tokens during the attention process.
-These additional vectors function similarly to support vectors in traditional kernel methods, providing additional decision boundaries within the self-attention mechanism.
-The introduction of these learnable support vectors not only mitigates the issue of insufficient tokens but also offers several other benefits.
-Firstly, they expand the learning capacity of transformers, allowing them to capture more nuanced patterns and relationships within the data.
-Secondly, these vectors stabilize the training process by ensuring that the model consistently attends to a fixed number of tokens, thereby maintaining the robustness and reliability of the attention mechanism across varying input sizes.
-Mixture of Attention Experts.
-While global attention mechanisms, such as the one we propose, offer significant advantages in terms of parallelizability and runtime efficiency, they are not without limitations.
-One key drawback is their restricted local receptive field, meaning that each token can only attend to the same subset of selected tokens within each layer.
-This limitation arises from the inherent design of global attention, where every token in the sequence attends to the same sampled tokens across all attention heads.
-However, upon closer examination, we realized that each attention head operates independently of the others.
-This independence implies that the outputs of the self-attention mechanism remain unchanged regardless of the token permutations applied within each attention head.
-To overcome this limitation, we drew inspiration from the Mixture of Experts (MoE) literature, particularly the concept of expert-choosing-token (Zhou et al., 2022).
-Our method introduces a novel approach where each attention head selects a different subset of sampled tokens to attend to.
-This process can be executed in parallel, as the results of the h-sampling without replacement are independent of one another.
-By enabling each attention head to focus on a distinct subset of tokens, our transformer architecture effectively increases the local receptive field by a factor of h, where h represents the number of attention heads.
-This means that with 4 attention heads and a 25% sampling rate, our model can approximate a full self-attention pattern when necessary, significantly enhancing its ability to capture both global and local dependencies within the data.
-Putting It All Together.
-Our transformer architecture integrates several cutting-edge techniques to maximize performance and efficiency.
-We adopt the Pre-RMSNorm formulation (Chen et al., 2018; Zhang and Sennrich, 2019), which has been proven to be both faster and more conducive to a higher convergence rate during training.
-This formulation normalizes the input before applying the self-attention mechanism, leading to improved stability and efficiency.
-For the residual branch of our model, we employ a zero-initialized learnable scalar multiplier (Bachlechner et al., 2021), a technique known to enhance the training dynamics of deep networks.
-However, we introduce a slight modification: the learnable scalar multiplier is shared across all layers of the network, promoting consistency and reducing the number of parameters that need to be learned.
-For importance score computation, a multi-layer perceptron transforming tokens’ latent is sufficient because the latents are complexly transformed by multiple transformer layers before.
-Modifications to integrate our sampling mechanism is described in earlier text.
-This integration allows our model to benefit from the efficiency and enhanced learning capacity provided by the differentiable top-k sampling and learnable support vector paddings.
-Finally, to align our model with the latest advancements in language technology, we incorporate flash attention (Dao et al., 2022; Dao, 2024).
-Flash attention is a highly optimized variant of self-attention that offers higher throughput and a lower memory footprint, making it ideal for large-scale models like ours.
-By combining these state-of-the-art techniques, our transformer architecture is not only capable of achieving top-tier performance across a variety of tasks but also does so with greater efficiency, ensuring it can be deployed effectively in real-world applications where computational resources and time are critical factors.
-4.4. Complexity Analysis
-Consider the input token $\mathbb{X}\in\mathbb{R}^{n\times d}$, query and key weights $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$, and $k$ sampled tokens, the computational cost for each 1-head transformer layer is as follows:
-  • Linear transformations for $\mathbb{Q},\mathbb{K}$ and $\mathbb{V}$: $O(nd^{2})$,
-  • Importance score of tokens for each attention head: $O(nd)$,
-  • Selecting top-$k$ importance score: $O(n+k)$ forward pass, $O(n+k^{2})$ backward pass with the formulation in Section 4.2,
-  • Attention from between $n$ queries and top-$k$ most important keys: $O(nkd_{h})$,
+Simply attending all query vectors to a set of sampled key and value vectors has a limitation: when the number of tokens is less than k (the number of sampled tokens), our algorithm fails.
+To address this issue, we introduce 2k learnable key and value vectors, each associated with learnable importance scores, which are sampled concurrently with input tokens.
+The process can be described mathematically as follows:
+$\mathbf{P}$ $\leftarrow\operatorname{Concatenate}(\mathbf{P},\mathbf{P_{\text{supp}}}),$
+$\mathbf{Z}$ $\leftarrow\operatorname{Concatenate}(\mathbf{Z},\mathbf{Z}_{\text{supp}})$
+where $\mathbf{P}:\mathbb{R}^{n\times 2d}$ is the key-value matrix, $\mathbf{Z}:\mathbb{R}^{n\times h}$ is the multi-head importance score matrix, $\mathbf{P_{\text{supp}}:\mathbb{R}^{2k\times 2d}}$ is the learnable support vector matrix, and $\mathbf{Z_{\text{supp}}:\mathbb{R}^{2n\times h}}$ is the learnable multi-head importance score matrix.
+These vectors and learned importance scores function is much similar to support vectors and Lagrange multipliers in traditional kernel methods, effectively introducing additional decision boundaries within the self-attention mechanism.
+Consequently, this not only enhances the learning capacity of transformers but also ensures stable training by consistently attending to a fixed number of tokens.
+Architecture Details.
+Our transformer architecture employs the Pre-RMSNorm formulation (Chen et al., 2018; Zhang & Sennrich, 2019), which has been demonstrated to accelerate convergence rates.
+For the residual branch, we utilize a zero-initialized learnable scalar multiplier (Bachlechner et al., 2021), with a modification: the learnable scalar multiplier is shared across all network layers.
+The self-attention modules in our architecture are adapted to incorporate our sampling mechanism, as discussed in detail.
+The importance score neural network for each self-attention module is implemented as a simple multi-layer perceptron, as the latent representations produced by the preceding transformer stack already exhibit sufficient complexity and expressiveness.
+Finally, to align with recent advancements in language modeling, we integrate flash attention (Dao et al., 2022; Dao, 2024) into our architecture, which offers higher throughput and reduced memory usage.
+The full algorithm is summarized in Algorithm 1, which is in the overview section.
+4.5 Complexity Analysis
+For an input token $\mathbb{X}\in\mathbb{R}^{n\times d}$ with query and key weights $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$ and $k$ sampled tokens, the computational costs per 1-head transformer layer are:
+  • Linear transformations for $\mathbb{Q},\mathbb{K},\mathbb{V}$: $O(nd^{2})$,
+  • Calculating token importance scores: $O(nd)$,
+  • Selecting top-k tokens: $O(n+k)$ forward, $O(n+k^{2})$ backward,
+  • Attention between $n$ queries and top-k keys: $O(nkd_{h})$,
   • FFN network: $O(nd^{2})$.
-Given these complexities and considering the long-ranged data structures which can implies $n\gg d$, the asymptotic computational complexity of one single SAMSA layer is $O(n)$.
-The asymptotic space complexity of one single SAMSA layer is also $O(n)$.
-5. Experiments
-Table 1.
-Performance/Runtime of SAMSA compared to the full self-attention (Vaswani et al., 2017) with different sampling sizes.
-Green and Red indicate better and worse performances/runtimes respectively.
-SAMSA Full Attention
-Sampling Size 128 128 256 256 512 512 N/A
+For large number of tokens ($n\gg d$), the asymptotic time and space complexity of SAMSA layer is $O(n)$.
+5 Experiments
+5.1 Experiment Design
+Our SAMSA models are either using hard/soft sampling indicating whether the straight through estimator is used or not.
+We have trained 45 models to measure the effectiveness of our method, both the performance and efficiency-wise of our model, against full self-attention (Flash Attention v2 (Dao, 2024)) and other specialized SOTA models on three data modalities (Sequence, Point Cloud, and Graph) on 4 datasets (Long Range Arena (all five), ModelNet40, ShapeNetPart, and Long Range Graph Benchmark (only Peptides)).
+To minimize any possible bias when comparing between baselines, we use no data augmentation and identical hyperparameters across different tasks.
+This also implies we conducted little hyperparameter tuning and mostly focused on the architectural design.
+5.2 Performance-Efficiency Trade-offs
+Table 1: Performance/Runtime of the full self-attention (Vaswani et al., 2017) compared to SAMSA with three different sampling sizes and two sampling formulations.
+Green and Red indicate SAMSA having better or worse statistics compared to full self-attention, respectively.
+The results of full self-attention in LRA are obtained from (Ma et al., 2023b).
+It has better performance than what reported in the dataset paper (Tay et al., 2021).
+The results of full self-attention in LRGB are obtained from the dataset paper (Dwivedi et al., 2022).
+The results of full self-attention on Shape Classification on ModelNet40 is run by ourselves, with the exact same hyperparameters as other SAMSA models.
+The results of PartSegmentation on ShapeNetPart is obtained from PointBERT (Yu et al., 2022).
+SAMSA Full Self-Attention
+Exp Settings Sampling Size 128 256 512 N/A
 Soft/Hard Sampling Soft Hard Soft Hard Soft Hard
-Sequence ListOPS 41.53 41.13 41.89 40.57 41.33 40.88 37.11
-Pathfinder 81.97 71.83 80.55 69.71 79.74 72.28 71.83
-Image 48.64 48.24 48.73 47.38 48.24 46.75 42.94
-Text 64.74 65.42 65.53 65.42 65.38 65.42 65.21
-Retrieval - 82.05 - 81.89 - 81.81 79.14
-Graph Pept-func 72.01 72.01 71.42 72.21 - 72.45 63.26
-Pept-struct 0.2519 0.2458 - 0.2515 - 0.2468 0.2529
-Point Cloud ModelNet40 91.11 90.91 91.72 90.95 91.35 91.39 91.07
-ShapeNetPart - 85.74 - 85.87 - 85.62 85.1
-Inference Speed (1024 tokens) 1.63×1.63$\times$1.63 × 2.17×2.17$\times$2.17 × 1.44×1.$4$ 1.62×1.62$\times$1.62 × 1.08×1.08$\times$1.08 × 1.44×1.$4$ 1$\times$
-5.1. Experimental Design
-Our SAMSA models utilize either hard or soft sampling techniques, determined by whether the reparameterization trick is employed or not.
-Soft sampling foregoes the reparameterization trick, instead operating with a very small Gumbel temperature (0.1), denoted as $g=g_{\text{soft}}$.
-This approach allows for a smoother, more probabilistic selection of tokens during training.
-Conversely, hard sampling employs the reparameterization trick with a standard Gumbel temperature (1.0), resulting in a more deterministic selection process.
-To rigorously evaluate the effectiveness of our approach, both in terms of performance and efficiency, we trained a total of 45 models.
-These models were benchmarked against full self-attention (specifically, Flash Attention v2 (Dao, 2024)) as well as other specialized state-of-the-art (SOTA) models across those three distinct data modalities:
-  • Sequence (multilinear)
-    • Long Range Arena (Tay et al., 2021)
-  • Point Cloud (non-multilinear)
-    • ModelNet40 (Wu et al., 2015)
-    • ShapeNetPart (Chang et al., 2015)
-  • Graph (non-multilinear)
-    • Long Range Graph Benchmark (Peptides) (Dwivedi et al., 2022; Singh et al., 2015)
-The selection of these datasets was intentional and rooted in a clear rationale.
-Our model, being a global attention model, was evaluated on datasets that highlight both the strengths and weaknesses typically associated with global attention mechanisms.
-Global attention models, while powerful in capturing long-range dependencies, are often criticized for their weakness in handling local structures.
-Several of the datasets we selected, such as 3 LRA datasets (Nangia and Bowman, 2018; Linsley et al., 2018; Krizhevsky and Hinton, 2009) and 2 LRGB datasets (Singh et al., 2015), are known to exhibit strong locality characteristics, making them ideal for testing this potential weakness.
-On the other hand, global attention models excel in tasks that require long-range modeling, a strength we tested using datasets that demand a global perspective, including 3 LRA datasets (Linsley et al., 2018; Maas et al., 2011; Radev et al., 2009), and 2 Point Cloud datasets (Wu et al., 2015; Chang et al., 2015).
-Furthermore, the chosen data modalities—sequence, point cloud, and graph—are fundamentally different from each other, with minimal shared inductive biases.
-This diversity ensures that our model’s generalizability is rigorously tested, as it cannot rely on the exploitation of inherent structures unique to any one modality.
-To ensure the validity and fairness of our comparisons with baseline models, we took several precautions to minimize bias.
-We avoided the use of data augmentation, which could introduce variability, and maintained nearly identical hyperparameters across different tasks.
-This consistent approach implies that we engaged in minimal hyperparameter tuning, allowing for a more accurate assessment of the SAMSA model’s inherent capabilities and its performance relative to other models under similar conditions.
-By controlling these variables, we aimed to present a clear and unbiased evaluation of SAMSA’s effectiveness across diverse data modalities and against a range of competitive benchmarks.
-Table 2. (Long Range Arena) Accuracy on the full suite of long range arena (LRA) tasks.
-Method ListOps Text Retrieval Image Pathfinder
-BigBird (Zaheer et al., 2020) 36.05 64.02 59.29 40.83 74.87
-Reformer (Kitaev et al., 2020) 37.27 56.10 53.40 38.07 68.50
-Performer (Choromanski et al., 2021) 18.01 65.40 53.82 42.77 77.05
-Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34
-Luna-256 (Ma et al., 2021) 37.98 65.78 79.56 47.86 78.55
-SAMSA-hard (Ours) 41.12 65.42 81.89 48.24 72.27
-SAMSA-soft (Ours) 41.88 65.53 82.05 48.73 81.97
-5.2. Sampling vs. Full-Self Attention
-We evaluated the performance of our SAMSA model across different sampling rates, with detailed results presented in Table 1.
-The experiments were conducted on nine different datasets using six distinct SAMSA model configurations.
-The results consistently demonstrate that SAMSA not only outperforms the full self-attention mechanism in most cases but also exhibits superior efficiency in both training and inference phases.
-This performance boost is particularly notable given the reduced computational overhead associated with SAMSA.
-Table 3. Experimental results on Peptides func and Peptides struct. We report the means and standard deviations of 4 random seeds.
+Sequence ListOPS L=2048 Acc 41.53 41.13 41.89 40.57 41.33 40.88 37.11
+Pathfinder L=1024 Acc 81.97 71.83 80.55 69.71 79.74 72.28 71.83
+Image L=1024 Acc 48.64 48.24 48.73 47.38 48.24 46.75 42.94
+Text L=2048 Acc 64.74 65.42 65.53 65.42 65.38 65.42 65.21
+Retrieval L=4096 Acc - 82.05 - 81.89 - 81.81 79.14
+Graph Pept-func L=450 AP 72.01 72.01 71.42 72.21 - 72.45 63.26
+Pept-struct L=450 MAE 0.2519 0.2458 - 0.2515 - 0.2468 0.2529
+Point Cloud ModelNet40 L=1024 Acc 91.11 90.91 91.72 90.95 91.35 91.39 91.07
+ShapeNetPart L=2500 IoU - 85.74 - 85.87 - 85.62 85.1
+Inference Speed (1024 tokens) $\mathbf{1.63\times}$ $\mathbf{2.17\times}$ $\mathbf{1.44\times}$ $\mathbf{1.62\times}$ $\mathbf{1.08\times}$ $\mathbf{1.44\times}$ 1$\times$
+We evaluated the performance of our SAMSA model on nine different datasets using six distinct SAMSA model configurations. The detailed results is presented in Table 1.
+Trade-offs against Full Self-Attention.
+At a 50% sampling rate, SAMSA is significantly faster and cheaper to compute than full self-attention, particularly due to the hard formulation, which avoids the need for gradient estimation during inference.
+Performance-wise, SAMSA consistently outperforms full self-attention, even at very low sampling rates (e.g., Retrieval: 128 / 4k).
+This superior performance is likely due to the regularization effect introduced by the sampling operator.
+Early in training, the random selection of tokens acts like DropAttention regularization (Zehui et al., 2019), preventing overfitting by ensuring attention is not overly concentrated on specific tokens.
+As training progresses, the sampler becomes more effective at identifying important tokens, reducing the regularization effect and allowing for more targeted attention.
+This dynamic aligns with known phenomena where early-stage regularization accelerates convergence, leading to improved overall performance (Liu et al., 2023b).
+Trade-offs between Different Sampling Rates.
+Efficiency-wise, Table 1 shows that SAMSA’s inference speed decreases with lower sampling rates and can be twice as fast as full self-attention on an A100.
+Performance-wise, the table also shows that a sampling rate of 256 tokens per head outperforms both full self-attention and higher sampling rates, suggesting that not all tokens equally impact attention effectiveness.
+This "sampling nonlinearity" may help filter out redundant tokens.
+Additionally, SAMSA’s performance on easy tasks like LRA-Text and LRA-Image decreases as sampling rates increase, indicating a potentially stronger regularization effect of lower sampling rates.
+Trade-offs between Hard and Soft Sampling.
+Hard-SAMSA is much faster than soft-SAMSA, with a 33% speed advantage at a 50% sampling rate.
+However, hard-SAMSA underperforms in certain tasks, likely due to discrepancies between the forward and backward passes in Gumbel-Softmax (Jang et al., 2017).
+Notably, hard-SAMSA’s learning curves in the Pathfinder task are volatile, reflecting the adverse effects of the straight-through estimator.
+Interestingly, earlier navigation through "breakpoints" leads to worse performance, possibly due to flat local minima, as shown by the stair-like learning curves (figure in Appendix E).
+However, it should be noted that hard-SAMSA performance is not inferior to soft-SAMSA with a counter-example: superior performance on graph data modality.
+5.3 Sampling vs. Specialized Models
+Aside from comparing against full self-attention, we also compare our model against the SOTA-specialized model on each of our experimented data modalities (sequences, graph, and point cloud).
+Sequences.
+Table 2: (Long Range Arena) Accuracy on the full suite of long range arena (LRA) tasks. Bold indicates best sequence-models while underline indicates best data modality agnostic models.
+Method ListOps Text Retrieval Image Pathfinder Modality
+BigBird (Zaheer et al., 2020) 36.05 64.02 59.29 40.83 74.87 Sequence
+Reformer (Kitaev et al., 2020) 37.27 56.10 53.40 38.07 68.50 Any
+Performer (Choromanski et al., 2021) 18.01 65.40 53.82 42.77 77.05 Any
+Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34 Any
+Luna-256 (Ma et al., 2021) 37.98 65.78 79.56 47.86 78.55 Any
+Transformer (Vaswani et al., 2017) 37.11 65.21 79.14 42.44 71.83 Any
+S4 (Gu et al., 2022) 88.65 76.02 87.09 86.09 86.05 Sequence
+MEGA (Ma et al., 2023b) 63.14 90.43 91.25 90.44 96.01 Sequence
+hard-SAMSA (Ours) 41.12 65.42 81.89 48.24 72.27 Any
+soft-SAMSA (Ours) 41.88 65.53 82.05 48.73 81.97 Any
+In Table 2, our model has competitive results against many other efficient transformers and the full attention transformer.
+From the table, it can be seen that MEGA (Ma et al., 2023b) and S4 (Gu et al., 2022) have the highest performance.
+This is true and expected because recurrent and locality via EMA are very strong inductive bias, making the model only work for sequences.
+Therefore, it is natural that our method, which is designed for multiple-data modalities, cannot outperform methods like those.
+Against weaker to no inductive bias like the rest, we achieved better performance on every task except text classification (losing Luna-256 by 0.25%).
+Graph.
+Table 3: Experimental results on Peptides func and Peptides struct.
+The results of methods other than ours are taken from their respective works and LRGB (Dwivedi et al., 2022).
+We report the means and standard deviations of 4 random seeds.
 Method
@@
 AP
-GCN (Kipf and Welling, 2017) 0.3496 ± 0.0013 0.5930 ± 0.0023
+GCN (Kipf & Welling, 2017) 0.3496 ± 0.0013 0.5930 ± 0.0023
 GINE (Xu et al., 2019) 0.6346 ± 0.0071 0.5498 ± 0.0079
-GatedGCN (Bresson and Laurent, 2017) 0.3420 ± 0.0013 0.5864 ± 0.0077
-GatedGCN + RWSE (Bresson and Laurent, 2017) 0.3357 ± 0.0006 0.6069 ± 0.0035
+GatedGCN (Bresson & Laurent, 2017) 0.3420 ± 0.0013 0.5864 ± 0.0077
+GatedGCN + RWSE (Bresson & Laurent, 2017) 0.3357 ± 0.0006 0.6069 ± 0.0035
 Drew-GCN + LapPE (Gutteridge et al., 2023) 0.2536 ± 0.0015 0.7150 ± 0.0044
 GraphGPS + LapPE (Rampášek et al., 2023) 0.2500 ± 0.0005 0.6535 ± 0.0041
-GRIT (Ma et al., 2023) 0.2460 ± 0.0012 0.6988 ± 0.0082
+GRIT (Ma et al., 2023a) 0.2460 ± 0.0012 0.6988 ± 0.0082
 Graph VIT (He et al., 2023) 0.2449 ± 0.0016 0.6942 ± 0.0075
@@
 MPNN + VN + RWSE (Cai et al., 2023) 0.2529 ± 0.0009 0.6685 ± 0.0062
-SAMSA-hard (ours) 0.2458 ± 0.0013 0.7221 ± 0.0042
-SAMSA-soft (ours) 0.2519 ± 0.0011 0.7171 ± 0.0051
-We hypothesize that SAMSA’s superior performance can be attributed to the regularization effect introduced by the sampling operator.
-During the early stages of training, the sampling mechanism tends to select tokens almost randomly, which is akin to the DropAttention approach (Zehui et al., 2019).
-This random selection acts as a form of regularization, preventing the model from overfitting by ensuring that attention is not overly focused on any particular subset of tokens.
-As training progresses and the sampler becomes more adept at identifying important tokens, the attention mechanism becomes more targeted, gradually diminishing the regularization effect.
-This dynamic behavior aligns with previously observed phenomena, where early-stage regularization accelerates the model’s convergence rate, leading to better overall performance (Liu et al., 2023b).
-Regarding the trade-off between performance and efficiency, the data in Table 1 reveals a significant insight: SAMSA can achieve superior results with a sampling rate of just 256 tokens per head, outperforming both full self-attention and configurations with higher sampling rates.
-This finding strongly suggests that not all tokens contribute equally to the attention mechanism’s effectiveness.
-Instead, there exists an optimal or ”Goldilocks-zone” sampling rate where the regularization effect of sampling maximizes performance while simultaneously reducing inference time.
-This balance between efficiency and effectiveness is a key advantage of SAMSA, as it highlights the potential to streamline attention computation without sacrificing accuracy.
-Table 4. Experimental results on ModelNet40 dataset (Wu et al., 2015) and part segmentation results on ShapeNetPart dataset (Chang et al., 2015) against point-cloud specialized baselines.
+24-layer hard-SAMSA (ours) 0.2458 ± 0.0013 0.7221 ± 0.0042
+24-layer soft-SAMSA (ours) 0.2519 ± 0.0011 0.7171 ± 0.0051
+In Table 3, our model has outperformed most methods in the two tasks we tested.
+This shows the efficacy of our newly proposed Graph-PE.
+Note that, our method’s learnable positional encoding is computed cheaply, and how cheap the computation is best represented in Table 4.
+The table shows that our entire model computes everything even before other methods finish their positional encoding computation.
+It also shows the efficacy of the SAMSA model as a very deep network, which is also different from traditional graph learning methods which cannot extend beyond 8 layers due to over-smoothing.
+Table 4: Runtime of Graph Positional Encoding on one epoch on A100 compared to our 24-layer SAMSA model. The runtime of LapPE and RWPE are taken from (Dwivedi et al., 2022).
+Method LapPE RWPE SAMSA (whole model)
+Peptides Struct 73s 53s 35s
+Table 5: Inference time (in seconds) of the networks for ModelNet40 classification test split in 1 A100 and 8 CPUs with a batch size of 32.
+The results of PointNet and DGCNN are taken from (Kaba et al., 2023).
+Method PointNet DGCNN hard-SAMSA (ours) soft-SAMSA (ours)
+Inference Time 18s 23s <1s 1s
+Table 6: Experimental results on ModelNet40 dataset (Wu et al., 2015) and part segmentation results on ShapeNetPart dataset (Chang et al., 2015) against point-cloud specialized baselines.
 Method ModelNet40 ShapeNetPart
@@
 Point Transformer (Zhao et al., 2021) 90.6 93.7 83.7 86.6
-SAMSA-hard (Ours) 88.9 91.3 83.8 85.8
-SAMSA-soft (Ours) 88.7 91.7 - -
-5.3. Sampling vs. Specialized Models
-A significant advantage of SAMSA is its applicability across a wide range of data modalities, making it a versatile tool in diverse fields.
-This versatility is evident in the performance comparisons shown in Table 2, Table 3, and Table 4, where SAMSA is evaluated against a variety of specialized methods in graph processing (using the Long Range Graph Benchmark) and point cloud analysis (using ModelNet40 and ShapeNetPart).
-Despite SAMSA being designed without any inherent inductive biases tailored to specific data types, it demonstrates a performance that is not only comparable to these specialized methods but also surpasses many graph-based models.
-In the context of graph processing, SAMSA produced competitive results on both the Peptides Func and Struct datasets.
-Notably, it achieved an average precision (AP) of $72.21\%$, outperforming all other graph methods evaluated in these benchmarks.
-This highlights SAMSA’s ability to effectively capture and utilize the important features within graph data, even when competing against models specifically designed for such tasks.
-For point cloud analysis, while SAMSA did not reach the same level of performance as in graph processing, primarily due to the absence of local information aggregation mechanisms which are crucial in this domain, it still delivered results that were on par with many models specifically optimized for point clouds.
-This demonstrates SAMSA’s robustness and adaptability, as it remains competitive even in scenarios where it faces inherent challenges.
-Overall, SAMSA’s ability to perform well across different data modalities underscores its potential as a foundational model, capable of replacing or complementing more specialized approaches in various applications.
-6. Conclusion
-In conclusion, the SAMSA (SAMpling-based Self-Attention) mechanism addresses the limitations of traditional transformers by reducing the quadratic complexity typically associated with self-attention.
+hard-SAMSA (Ours) 88.9 91.3 83.8 85.8
+soft-SAMSA (Ours) 88.7 91.7 - -
+Point Cloud.
+Performance-wise, Table 6 shows that SAMSA has competitive results against other point cloud models.
+This is mainly because our model does not include any data-modality-specific inductive bias (like the usage of kNN for locality heuristic); therefore, it is susceptible to overfitting in scenarios where data is not abundant.
+Our method is also much faster than others, up to 18$\times$ faster than PointNet (which is the most small, simple, efficient point cloud model), as shown in Table 5.
+6 Conclusion
+SAMSA (SAMpling-based Self-Attention) mechanism addresses the limitations of traditional transformers by reducing the quadratic complexity typically associated with self-attention.
 Unlike other efficient transformers that are often data-modality-specific, SAMSA provides a versatile, context-aware linear complexity solution that can be applied across various data modalities.
@@
 Limitations.
-Despite the promising results achieved with SAMSA, there are certain limitations on the biased gradients of hard sampling formulation, which makes our model hard to converge in more difficult problems (Pathfinder).
-The biased gradient issue in hard sampling methods is mostly from the misestimation of gradient of the non-chosen path.
-We believe that future work on more precise gradient estimation would yield inference-fast model with much better accuracy.
+Despite the promising results achieved with SAMSA, there are certain limitations on the misaligned gradients of hard sampling formulation, which makes our model hard to converge in more difficult optimization problems (Pathfinder).
+This misaligned gradient issue in hard sampling methods is mostly from the misestimation of gradients of the non-chosen path.
+We believe that future work should focus on more precise gradient estimation would result in an inference-fast model with much better accuracy.
 Appendix A Datasets
@@
 Long Range Arena.
-The Long Range Arena (LRA) (Tay et al., 2021) is a composition benchmark of 5 tasks: ListOps (Nangia and Bowman, 2018), ImDB review (Maas et al., 2011), ACL Anthology Network (Radev et al., 2009), Grayscaled CIFAR-10 (Krizhevsky and Hinton, 2009), and Pathfinder (Linsley et al., 2018).
+The Long Range Arena (LRA) (Tay et al., 2021) is a composition benchmark of 5 tasks: ListOps (Nangia & Bowman, 2018), ImDB review (Maas et al., 2011), ACL Anthology Network (Radev et al., 2009), Grayscaled CIFAR-10 (Krizhevsky & Hinton, 2009), and Pathfinder (Linsley et al., 2018).
 These five tasks are all classification and they feature very long sequences: ListOps (2,048 tokens), ImDB review (1,024 tokens), ACL Anthology Network (4,096 tokens), Grayscaled CIFAR-10 (1,024 tokens), and Pathfinder (1,024 tokens).
@@
 We use the pytorch framework (Paszke et al., 2019) to conduct our empirical experiments.
-Here is the list of hyperparameters:
-  • Optimizer: AdamW (Loshchilov and Hutter, 2019)
+Here is the list of the common hyperparameters:
+  • Optimizer: AdamW (Loshchilov & Hutter, 2019)
+  • Batch size: 32
   • Learning rate: 0.001
   • $\beta_{1},\beta_{2}$: 0.9, 0.999
-  • Weight decay: 0.01
   • Learning scheduler: Cosine Annealing
-  • Learning rate warm-up: Untuned Linear Warmup (Ma and Yarats, 2021)
+  • Learning rate warm-up: Linear Warmup-2000 steps
   • Gradient Clipping: Max Gradient Norm (2.0)
-Table 5. Training details.
-Task Dataset Batch Size No. of Epochs
-Point Cloud ModelNet40 32 600
-ShapeNetPart 32 600
-Sequence ListOPS 32 50
-Text 32 50
-Retrieval 32 50
-Image 32 50
-Pathfinder 32 50
-Graph Peptides-struct 32 240
-Peptides-func 32 160
 Our training phase details can be visited in Table 7.
+Table 7: Training hyperparameter details.
+Task Dataset Learning Rate Gradient Clipping Weight Decay No. of Epochs
+Point Cloud ModelNet40 0.002 2.0 0.1 200/600
+ShapeNetPart 0.002 2.0 0.1 600
+Sequence ListOPS 0.001 2.0 0.01 50
+Text 0.001 2.0 0.01 50
+Retrieval 0.001 2.0 0.01 50
+Image 0.001 2.0 0.01 50
+Pathfinder 0.001 2.0 0.01 50
+Graph Peptides-struct 0.001 2.0 0.01 240
+Peptides-func 0.0004 2.0 0.01 160
 As for the input preprocessing, our works are as follows:
@@
   • Long Range Graph Benchmark: We input the whole graph, with both node and edge as tokens.
-  • ModelNet40: Following previous work (Charles et al., 2017), we use farthest point sampling to sample 1024 initial points.
+  • ModelNet40: Following previous work (Charles et al., 2017), we use farthest point sampling to sample 1024 initial points with normals.
   • ShapeNetPart: We operate on the initial 2,500 point-normal without any change to the dataset.
-All of the training processes utilize Cross Entropy Loss, without label smoothing normalization.
+[…]
+Appendix C Architectural Specification
+Table 8 shows our model’s architectural hyperparameters, for reproducibility purposes.
+Table 8: Architectural details.
+Dataset $\mathbf{n_{\textbf{depth}}}$ $\mathbf{d_{\textbf{model}}}$ $\mathbf{d_{\textbf{FFN}}}$ $\mathbf{n_{\textbf{heads}}}$ $\mathbf{p_{\textbf{dropout}}}$ $\mathbf{p_{\textbf{droppath}}}$
+ModelNet40 6 256 768 8 0.1 0.1
+ShapeNetPart 8 256 768 8 0.1 0.1
+LRA-ListOPS 6 256 768 8 0.1 0.0
+LRA-Pathfinder 6 128 512 8 0.3 0.0
+LRA-Text 4 128 512 8 0.3 0.0
+LRA-Image 8 128 512 8 0.3 0.0
+LRA-Retrieval 6 128 512 8 0.3 0.0
+LRGB-Peptides-Func 24 128 512 8 0.3 0.2
+LRGB-Peptides-Struct 24 128 512 8 0.3 0.2
+LRA Tasks use Sinusoid-1D positional encodings.
+Appendix D Miscellaneous
+ST-Gumbel-Sigmoid’s Formula:
+$\operatorname{GumbelSigmoid}(\pi)$ $\triangleq\frac{1}{1+\exp\left((\pi_{j}+g_{j})/\tau\right)},$
+$\operatorname{ST-GumbelSigmoid}(\pi)$ $\triangleq 1\begin{cases}1,&\text{if }\pi\leq 0,\\ 0,&\text{otherwise},\end{cases}$
+$\partial{\operatorname{ST-GumbelSigmoid}}/\partial{\pi}$ $\triangleq\partial{\operatorname{GumbelSigmoid}}/\partial{\pi},$
+where $g_{i}$ are independent and identically samples from Gumbel(0,1) distribution, $\pi_{i}$ are the unnormalized logits, and $\tau$ is the temperature parameter that controls the smoothness of the distribution.
+Appendix E Learning Curves of Our Models
+Figure 2,  3,  4 show the learning curve of SAMSA in various settings, providing for future understanding of how our model learned. All of the lines are plotted with EMA-5 smoothing.
+Figure 2: Learning Curves of SAMSA models in Sequence Tasks
+Figure 3: Learning Curves of SAMSA models in Graph Tasks
+Figure 4: Learning Curves of SAMSA models in Point Cloud Tasks